### v3

# FaithLM: Towards Faithful Explanations for Large Language Models

Abstract
Large Language Models (LLMs) have become proficient in addressing complex tasks by leveraging their extensive internal knowledge and reasoning capabilities. However, the black-box nature of these models complicates the task of explaining their decision-making processes. While recent advancements demonstrate the potential of leveraging LLMs to self-explain their predictions through natural language (NL) explanations, their explanations may not accurately reflect the LLMs’ decision-making process due to a lack of fidelity optimization on the derived explanations. Measuring the fidelity of NL explanations is a challenging issue, as it is difficult to manipulate the input context to mask the semantics of these explanations. To this end, we introduce FaithLM for explaining the decision of LLMs with NL explanations. Specifically, FaithLM designs a method for evaluating the fidelity of NL explanations by incorporating the contrary explanations to the query process. Moreover, FaithLM conducts an iterative process to improve the fidelity of derived explanations. Experiment results on three datasets from multiple domains demonstrate that FaithLM can significantly improve the fidelity of derived explanations, which also provides a better alignment with the ground-truth explanations.
1 Introduction
Large language models (LLMs) exhibit remarkable performance in various natural language processing tasks, such as the GPT4 [1], LLaMA [39], and Claude [4]. However, these language models are commonly regarded as intricate black-box systems. The opacity of their internal mechanisms poses a significant challenge when trying to explain their decision-making process. The lack of transparency in LLMs, especially in API-accessed LLM services, inferences contradict the practical requirements of stakeholders and are in opposition to regulatory standards in various domains, such as GDPR [15, 14]. The imperative arises to develop explainability mechanisms for LLMs, particularly for their use in high-stakes applications such as healthcare.
Numerous studies have attempted to enhance the transparency of decision-making processes in LLMs by providing natural language (NL) explanations. Advanced billion-level LLMs, such as GPT4 [1], mainly leverage their reasoning abilities and internal knowledge for complex downstream problems. However, this complexity poses challenges to faithfully explain the underlying reasons behind their decisions with natural language sentences. In this manner, recent advancements are struggling to generate reliable NL explanations for interpreting LLMs [45]. Some work attempt to leverage powerful LLMs [27, 9] with auxiliary information to generate NL sentences or chain-of-thought (CoT) reasoning [22, 34, 8] as the explanations to explain the decision-making process of LLMs. Although existing work emerged that LLMs may possess the ability to self-explain [26, 42], or follow chain-of-thoughts to explain [25], their explanation-generating process usually overlooks the fidelity, a fundamental metric for evaluating the quality of explanations [12, 41]. Thus, the derived NL explanations may not faithfully reflect the decision-making process of LLMs [46, 40]. Measuring the fidelity of NL explanations is a challenging issue, as the NL explanations may provide crucial information beyond the input context. This makes it hard to mask the semantics of NL explanations by manipulating the input tokens.
Question and LLM Answer Faithful NL Explanation Question conditioned with Contrary NL Explanation
Question: Can the positive pole from two magnets pull each other closer? Original Answer: No Each magnet has a positive pole and a negative pole, and similar poles push each other away. Question: Each magnet has a positive pole and a negative pole, and similar poles pull each other closer. Can the positive pole from two magnets pull each other closer? New Answer: Yes
Table 1: An example of measuring fidelity of NL explanations. The LLM first answers the question in No. Given a faithful explanation “similar poles pull each other away," with its contrary NL explanation, the LLM changes the answer from No to Yes when introduced contrary NL explanation as an extra condition to LLM. This example indicates that the contrary explanation interrupts the LLM’s original prediction process with faithful information, demonstrating that the information in an original explanation is faithful and aligns with LLM’s initial answer.
To overcome this challenge, we propose a method to measure the fidelity of NL explanations. We give an example to convey the motivation in Figure 1. Specifically, the fidelity of an explanation can be measured by leveraging its contrary explanation as extra conditions of the input context, and observing the LLM’s output difference compared with its initial output. Here, a contrary explanation refers to a statement with opposite semantics to the original explanation. By incorporating the contrary explanation to the input context, we can identify an explanation as high fidelity if there is a significant change in the LLM’s output, such as from No to Yes. This change indicates that the crucial information present in the original explanation is substituted with the opposite meaning context in the contrary explanation, where the crucial information is essential to the LLM’s decision-making process. Based on this observation, we propose to extend the applicability of fidelity to the evaluation on NL explanations. This extension follows the integration of contrary explanations to represent the concepts of masking important features in traditional fidelity measurement.
Building upon this approach for measuring the fidelity of NL explanations, we introduce Faithful LLM Explainers (FaithLM) to generate faithful NL explanations for LLMs. Specifically, FaithLM adopts LLMs as explainers to generate the NL explanations and explanation trigger prompts, and iteratively optimizes the derived NL explanations and trigger prompts with the goal of fidelity enhancement. During the iterative process, FaithLM computes the fidelity of each derived explanation and optimized prompt based on our proposed fidelity measurement method, and progressively improves their fidelity through in-context learning. To demonstrate the effectiveness of FaithLM, we conducted several experiments on four different LLMs under three datasets from different domains. FaithLM achieved significantly higher fidelity performance in generating NL explanations and more closely matched the golden explanations compared with state-of-the-art baseline methods. In summary, our contributions can be summarized as follows:
  • Fidelity of NL Explanations: We propose to measure the fidelity of NL explanation by incorporating its contrary explanation into the query and observing LLM output difference.
  • Faithful LLM Explainers: We introduce FaithLM for improving the fidelity of NL explanations, aiming at faithfully explaining the decision-making process of LLMs.
  • Fidelity and Truthfulness: Experimental results on three datasets show that FaithLM can effectively improve the fidelity of NL explanations, which also shows a better alignment with the ground-truth explanations than state-of-the-art baseline methods.
2 Preliminaries
2.1 Notations and Objectives
We aim to explain the decisions of arbitrary targeted LLMs $f(\cdot)$ with NL explanations in a post-hoc manner. Given an input $\boldsymbol{X}$, the targeted LLMs generate an output $\boldsymbol{Y}=f(\boldsymbol{X})$. Our objective is to produce an NL explanation $\mathcal{E}_{\text{NL}}$ that faithfully explains the reasons behind the prediction of $\boldsymbol{Y}=f(\boldsymbol{X})$. In this work, we employ an LLM as the explainer $\textsl{g}(\cdot)$ to generate the NL explanation $\mathcal{E}_{\text{NL}}=\textsl{g}(\cdot~{}|~{}\boldsymbol{X},\boldsymbol{Y})$. However, the directly generated $\mathcal{E}_{\text{NL}}$ under single-forward passing may not be faithful and accurate, degrading the user’s trust in the prediction made by the targeted LLM. To this end, our work focuses on benefiting the explainer $\textsl{g}(\cdot)$ to generate more faithful NL explanations regarding the decision of $f(\cdot)$ in a post-hoc manner, where $f(\cdot)$ can be either closed-source or open-source LLMs.
2.2 Post-hoc Explanations of LLMs
Recent research on LLM explanations has primarily focused on post-hoc approaches due to the inaccessibility of LLM weights and architectures. One group of studies calculates importance scores for specific tokens [24, 17], and the other line of work generates natural language sentences [21, 7, 9, 28] and CoT reasoning [25, 8] as explanations by leveraging LLMs. Despite the advanced capability of powerful LLMs to generate NL explanations, issues with unreliability and non-fidelity persist. While studies like Chen et al. [6] utilize internal model weights aiming for reliable explanations generation, accessing these weights is challenging with closed-source LLMs. Considering the non-accessibility of LLMs, such as closed-source LLMs, our goal is to faithfully generate NL explanations instead of exploring the internal mechanisms or neurons of LLMs to generate explanations.
2.3 Limitations of Traditional Fidelity on Natural Language Explanations
The fidelity metric measures the fidelity of the given explanation, which is broadly applicable when ground-truth explanations are unavailable. In the NLP scenario, fidelity has been used to evaluate the heatmap-formatted explanations [24, 17], where the heatmap one highlights the important tokens of the input. Specifically, fidelity evaluates the explanation by removing the important tokens from the input $\boldsymbol{X}$ and checking the prediction difference of the targeted LLM. Following the definition of fidelity [29]. Given a sequence of tokens $\boldsymbol{I}=\{t_{1},\cdots,t_{M}\}\subseteq\mathcal{E}_{\text{NL}}$, which is identified as an important component of explanation to the prediction of a targeted LLM $\boldsymbol{Y}=f(\boldsymbol{X})$. Following the traditional fidelity definition, the fidelity can be estimated as follows:
$\mathrm{Fidelity}=f(\boldsymbol{X})-f(\boldsymbol{X}\setminus\boldsymbol{I}),$
where "$\boldsymbol{X}\setminus\boldsymbol{I}$" denotes the token removal from $\boldsymbol{X}$ in $\boldsymbol{I}$.
If important component $\boldsymbol{T}$ achieves higher fidelity, this demonstrates that $\mathcal{E}_{\text{NL}}$ comprises the crucial tokens that significantly influence the predictions of the targeted LLMs. However, it is challenging to evaluate the fidelity of NL explanations throughout the fidelity defined above, as the critical components in NL explanations may not contain in the input context $\boldsymbol{X}$. For instance, as the example shown in Figure 1, the critical component “similar poles push each other away" in $\mathcal{E}_{\text{NL}}$ does not explicitly appear in the input context. Thus, we cannot simply remove critical components from the question following the traditional definition. Therefore, in this paper, we propose a solution to systematically address this obstacle by removing the critical components from the semantic level instead of the token level.
3 FaithLM: The Explainer LLM Framework
In this section, we systematically introduce the generative explanation framework, FaithLM, which derives faithful explanations in natural language. The derived explanations are expected to accurately reflect the predictive decision-making process of targeted LLMs with high fidelity. Specifically, we first introduce the Fidelity Evaluator in Section 3.1 to quantitatively assess the fidelity of the NL explanations. Subsequently, FaithLM iteratively optimizes the fidelity of the derived explanations and searches for improved explanation trigger prompts using the proposed Fidelity Evaluator. The optimization frameworks for these processes are detailed in Sections 3.2 and 3.3.
3.1 Fidelity Evaluator for Natural Language Explanations
Due to the limitation of the fidelity metric described in Section 2.3, we introduce the Fidelity Evaluator to directly assess the fidelity of NL explanations, as shown in Figure 1. Specifically, we first extend the original fidelity definition to evaluate the fidelity of NL explanations $\mathcal{E}_{\text{NL}}$. Finally, we introduce the fidelity evaluator that follows the proposed fidelity measurement criteria.
Fidelity of NL Explanations.
Figure 1: The framework of Fidelity Evaluator. The evaluator calculates the fidelity scores of the derived explanations based on its contrary explanations.
To assess the fidelity of NL explanations, we extend the traditional fidelity definition to equip it with special constraints regarding NL explanations. To address this challenge, we propose formulating the fidelity of NL explanation as the prediction difference caused by the “contrary explanation". Specifically, the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ is defined as a statement obtained opposite semantics to the given NL explanation $\mathcal{E}_{\text{NL}}$. For instance, if the explanation is "similar poles push each other away," then the contrary explanation would be "similar poles pull each other closer." To estimate the fidelity of $\mathcal{E}_{\text{NL}}$, we use the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ as an extra given condition to the input queries, forming conditional LLM inferences $f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$. This operation results in different prediction results for targeted LLMs compared to the original predictions, which compels the targeted LLMs to follow contrary information from the explanations of the input queries. In this manner, the fidelity $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$ can be estimated by the prediction difference:
$\mathcal{S}_{\mathcal{E}}:=f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}}).$ (1)
This formulation aligns with the traditional definition of fidelity by observing the prediction difference of the LLM when considering input with and without the component supplied by the explanations.
Fidelity Evaluator.
Unreliable explanations derived from LLMs often present incorrect predictions [45], which means that the relationship between the reliability of the explanations and the correctness of predictions is significantly high. Motivated by the observation, we propose a framework of the Fidelity Evaluator as illustrated in Figure 1. Specifically, we utilize a powerful LLM agent, such as GPT-3.5, to generate the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ regarding the original $\mathcal{E}_{\text{NL}}$. The generation process for $\neg\mathcal{E}_{\text{NL}}$ is guided by the prompt provided in Appendix G. Given the contrary explanation as a condition for the input context, the Fidelity Evaluator computes the fidelity scores according to the difference of output logits $f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$. Intuitively, if the targeted LLMs’ output probability changes significantly enough to flip the result (i.e., judgements or choices), it implies that the contrary explanations $\neg\mathcal{E}_{\text{NL}}$ consist of the contrasting key messages against $\mathcal{E}_{\text{NL}}$ that can significantly impact the decision-making process of the targeted LLM. This suggests that the explanation $\mathcal{E}_{\text{NL}}$ contains key components that significantly support the inference process of the targeted LLM.
3.2 FaithLM for Fidelity-enhanced Explanation
In this section, we introduce an iterative framework designed to progressively enhance the fidelity of NL explanations. The primary goal of FaithLM here is to generate faithful NL explanations with iterative fidelity-enhanced optimization.
Fidelity-enhanced Explanation.
The framework of Fidelity-enhanced Explanation is illustrated in Figure 2(a). Since the initial explanation may be unreliable and unfaithful, we propose a fidelity-enhanced optimization approach designed to progressively generate explanations with higher fidelity. We aim to explain the response $\boldsymbol{Y}$ produced by the targeted LLM $f(\cdot)$ in response to the given input queries $\boldsymbol{X}$ with NL explanations $\mathcal{E}_{\text{NL}}$ following the goal of fidelity enhancement. In the first round of enhancement, the LLM explainer generates NL explanations $\mathcal{E}_{\text{NL}}$ following human-crafted explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ provided in Appendix G. The explanations are then generated by the explainer $\textsl{g}(\mathcal{P}_{\mathcal{E}}|~{}\boldsymbol{X},\boldsymbol{Y})$. Starting from the second round till converge, FaithLM collects a trajectory $\mathcal{T}$ with the NL explanations $\mathcal{E}_{\text{NL}}$ and their corresponding fidelity scores $\mathcal{S}_{\mathcal{E}}$ generated by Fidelity Evaluator. The collection process can be represented as $\mathcal{T}\leftarrow\{\mathcal{T},~{}[\mathcal{E}_{\text{NL}},~{}\mathcal{S}_{\mathcal{E}}]\}$, where $\mathcal{T}$ initially starts as an empty trajectory. Following this trajectory, the LLM explainer generates new explanations with the goal of achieving higher fidelity scores in subsequent iterations. This fidelity enhancement process is guided by the system prompts detailed in Figure 13 of Appendix H.
The trajectory $\mathcal{T}$ is continuously updated by incorporating each newly derived explanation with its assessed fidelity score until the convergence. Regardless of any given explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$, FaithLM can all systematically guide the generation of NL explanations, progressively improving fidelity scores by following the reference path established in the trajectory.
(a) Fidelity-enhanced Explanation
(b) Trigger Prompt Optimization
Figure 2: An overview of FaithLM framework for two differenent optimization objectives. The blue dotted line reveals the trajectory to optimize the NL explanation (Section 3.2), and the red dotted line indicates the trajectory of the explanation trigger prompt optimization (Section 3.3). “Traj. Prompt" denotes the trajectory system prompt shown in Section H.
Algorithm 1 FaithLM for fidelity-enhanced explanation
Input: Input $\boldsymbol{X}$, output $\boldsymbol{Y}$, targeted LLMs $f(\cdot)$, human-crafted prompt $\mathcal{P}_{\mathcal{E}}$, and LLM explainer $\textsl{g}(\cdot)$.
Output: NL explanation $\mathcal{E}_{\text{NL}}$.
1:  $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}~{}|~{}\boldsymbol{X},\boldsymbol{Y})$
2:  $\mathcal{T}=\varnothing$
3:  while  steps not end and decision not flips  do
4:     Estimate the fidelity score $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$
5:     Append $\mathcal{T}\leftarrow~{}\mathcal{T}\cap[\mathcal{E}_{\text{NL}},~{}\mathcal{S}_{\mathcal{E}}]$
6:     Update $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{T}~{}|~{}\boldsymbol{X},\boldsymbol{Y})$
7:  end while
Algorithm of Fidelity-enhanced Explanation.
The outline of FaithLM for Fidelity-enhanced Explanation is detailed in Algorithm 1. Specifically, in the first iteration, FaithLM generates the NL explanations using the human-craft prompts (line 1). starting from the second iteration till the convergence or optimization ends, FaithLM estimates the fidelity of the derived NL explanations (line 4). Then, we incorporate the explanation and its corresponding fidelity score to the trajectory (line 5), and update the explanations with the goal of achieving higher fidelity scores in subsequent iterations (line 6). The iteration terminates at a predetermined step or ceases earlier as soon as FaithLM observes a flipping performance from the targeted LLM $f(\cdot)$.
3.3 FaithLM for Trigger Prompt Optimization
Despite the success of enhancing fidelity in Section 3.2, the low quality of the explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ may still hinder the optimization process of receiving a high-fidelity explanation. Given that the unknown preference for prompts from LLMs, human-crafted trigger prompts used in Fidelity-enhanced Explanation Optimization might lead to sub-optimal fidelity enhancement in the derived explanations. In this section, we hereby propose a new optimization pipeline under FaithLM, aiming to optimize the trigger prompt $\mathcal{P}_{\mathcal{E}}$ for generating explanations with higher fidelity.
Trigger Prompt Optimization.
The framework of Trigger Prompt Optimization is shown in Figure 2(b). The framework aims to optimize the trigger prompt to generate NL explanations with higher fidelity. Different from the optimization goal in Section 3.2, the trajectory in this task collects the trigger prompts $\mathcal{P}_{\mathcal{E}}$ and their fidelity scores $\mathcal{S}_{\mathcal{P}}$. The trajectory is constructed by the system optimization prompts detailed in Figure 12 of Appendix H.
To estimate the fidelity score for a trigger prompt, FaithLM first adopts the randomly human-crafted trigger prompt to guide the LLM explainers to generate NL explanations, and then utilize the Fidelity Evaluator to assess the fidelity of the derived explanation. The final estimated score is averaged by the fidelity score $\mathcal{S}_{\mathcal{E}_{i}}$ of each instance $i$ from the hold-out dataset $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\in\mathcal{D}$. Formally, the fidelity score for a trigger prompt $\mathcal{P}_{\mathcal{E}}$ can be estimated as follows:
$\mathcal{S}_{\mathcal{P}}=\mathbb{E}_{\mathcal{E}_{i}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})}\big{[}\mathcal{S}_{\mathcal{E}_{i}}\big{]},$ (2)
where $\mathcal{S}_{\mathcal{E}{i}}$ represents the fidelity score of the explanation $\mathcal{E}_{i}$, which is generated by $\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$, as assessed by the Fidelity Evaluator discussed in Section 3.1.
During the optimization procedure, the trajectory begins from an empty set and starts to incorporate newly derived trigger prompts along with the fidelity scores in each optimization iteration. Following this trajectory, the LLM explainer generates a new trigger prompt with the goal of achieving higher fidelity scores of explanations in subsequent iterations. After several rounds of iterations, FaithLM ultimately yields an optimal explanation trigger prompt with the highest fidelity score for the LLM explainer to generate a more faithful NL explanation.
Algorithm 2 Trigger Prompt Optimization.
Input: Hold-out dataset $\mathcal{D}$, Targeted LLMs $f(\cdot)$, and LLM explainers $\textsl{g}(\cdot)$.
Output: Optimal explanation trigger prompt $\mathcal{P}_{\mathcal{E}}$.
1:  Initialize human-crafted $\mathcal{P}_{\mathcal{E}}$
[…]
3:  while  (Steps Not End)  do
4:     for $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\sim\mathcal{D}$ do
5:        $\mathcal{E}_{i}\leftarrow\textsl{g}(\mathcal{P}_{\mathcal{E}}~{}|~{}\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$
6:        Estimate the fidelity score $\mathcal{S}_{i}$ of $\mathcal{E}_{i}$
7:     end for
8:     $\mathcal{S}_{\mathcal{P}}=\mathbb{E}_{\mathcal{E}_{i}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})}\big{[}\mathcal{S}_{\mathcal{E}_{i}}\big{]}$
9:     Append $\mathcal{T}\leftarrow~{}\mathcal{T}\cap(\mathcal{P}_{\mathcal{E}},\mathcal{S}_{\mathcal{P}})$
10:     Update $\mathcal{P}_{\mathcal{E}}\leftarrow\textsl{g}(\mathcal{P}_{\mathcal{E}}~{}|~{}\mathcal{D})$
11:  end while
Algorithm of Trigger Prompt Optimization.
The outline of FaithLM for Trigger Prompt Optimization is detailed in Algorithm 2, which focuses on optimizing the trigger prompt for generating NL explanations. Specifically, in each iteration, LLM explainer $\textsl{g}(\cdot)$ leverages the trigger prompt to generate the NL explanations and estimates its fidelity (lines 4-7). The fidelity scores of the trigger prompts average the fidelity scores of the entire hold-out dataset (lines 8). Afterward, the trajectory appends the trigger prompt with its corresponding fidelity score (line 9), and updates the trigger prompt as a new sequence of words to achieve higher fidelity scores (line 10). Through multiple iterations, FaithLM progressively guides the trigger prompt to generate explanations with higher fidelity scores, following the reference path established in the trajectory. The iteration process terminates at a predetermined step, and we set 20 in our experiments.
4 Experiment
In this section, we conduct experiments to evaluate the performance of FaithLM, aiming to answer the following three research questions: RQ1: How does FaithLM perform in generating explanations in terms of efficacy? RQ2: Can optimized explanation trigger prompts transfer between different datasets? RQ3: Does the configurations of LLMs affect the explanation performance of FaithLM?
4.1 Dataset and Baseline
Datasets. We evaluate FaithLM on three datasets with multiple tasks: ECQA [2] dataset on commonsense question-answer task, TrivaQA-Long [5, 19] dataset on reading comprehension task, and COPA [20, 35] dataset on commonsense causal reasoning task. More details of datasets are provided in Appendix A. Baseline Methods. We compare FaithLM with two state-of-the-art baseline methods: SelfExp [26] and Self-consistency [42]. The former ones instruct LLMs to generate explanations using prompt engineering under single-forward inference, and the later ones leverage the outputs from the chain-of-thought prompting process as the model explanations.
4.2 Experiment Settings
In this part, we introduce the experimental settings for evaluating FaithLM. Two distinct types of explanation tasks and evaluation settings are illustrated as follows.
Fidelity-enhanced Explanation
In this task, our goal is to produce NL explanations that exhibit a higher degree of fidelity. The fidelity is exploited as a metric to evaluate fidelity. FaithLM is evaluated across all testing instances, where an NL explanation is generated for each instance, and the averaged fidelity score is calculated, serving as the reported metric to evaluate fidelity.
Explanation Trigger Prompt Optimization.
In this task, we aim to optimize the explanation trigger prompt that benefits FaithLM in generating better explanations. The optimization process is conducted on the same dataset, where 30 instances are sampled as a hold-off dataset in each optimization step from the training set. During the optimization process, the fidelity score of a trigger prompt is calculated as the average of the fidelity scores from the selected instances.
Evaluation Metrics.
The quality of the derived NL explanation is evaluated under the fidelity and truthfulness metrics. The fidelity follows Section 3.1, which observes the flipping rate of the targeted LLMs by incorporating contrary explanations to the input. The evaluation of truthfulness assesses the correlation between the derived NL explanations to the ground-truth explanations Specifically, we leverage GPT-3.5 and two well-trained natural language inference (NLI) models, Roberta-Large and XLNet-Large [31] from the huggingface hub [43], as the evaluators. With the same evaluators setup, the truthfulness evaluation follows the settings from [23], and uses the evaluation prompt provided in Appendix F. Specifically, the evaluators assess the derived explanations and ground-truth explanations, determining whether the two sentences belong to “similar content", “dissimilar content," or “non-relevant content". The higher proportion of “similar content", the more consistent results with the ground-truth NL explanations.
Implementation Details.
In the experiments, we explore two variants of LLMs as the targeted LLMs $f(\cdot)$: Vicuna-7B [11] and Phi-2 [18], two types of LLMs as the explainers $\textsl{g}(\cdot)$ in FaithLM: GPT-3.5-Turbo and Claude-2 [3]. The LLM agent for generating the contrary explanations takes the same LLMs as those used by the explainers. All reported results are calculated from the average scores of 3 times repetitions. The hyper-parameter settings and device configuration, including temperature and total optimization steps, of FaithLM are given in Appendix C and D, respectively.
4.3 Explanation Efficacy of FaithLM (RQ1)
In this section, we explore the efficacy of FaithLM from two perspectives. Specifically, we first evaluate the efficacy of NL explanations generated by Algorithm 1, and then assess the explanation trigger prompts generated by Algorithm 2.
Efficacy of Derived Explanations.
We assess the efficacy of derived explanations under the fidelity metric. FaithLM adopts the trajectory system prompts in Figure 13 of Appendix H. The generation of contrary explanations is guided by the prompt in Table 8 of Appendix G.
Figure 3: The fidelity evaluation of explanations from FaithLM on ECQA dataset (left), TriviaQA-Long (middle), and COPA dataset (right). The reported scores are the average fidelity on testing instances in each step of fidelity-enhanced optimization.
Figure 4: Trustfulness evaluation of the NL explanations. The higher the proportion of “similar to ground-truth explanation" obtained from classifiers, the more consistent the derived explanations are with the ground-truth NL explanations.
  • Fidelity Evaluation. The results in Figure 3 demonstrate that FaithLM achieves significantly higher fidelity scores across all three datasets compared with two baselines after 20 steps of optimization. Moreover, the optimization curve of fidelity demonstrates that 20 rounds of optimization are sufficient to converge. Similar phenomenon occurs across different settings of explainers and targeted LLMs. Additional results are provided in Appendix E.
  • Truthfulness Evaluation. To evaluate the truthfulness of the derived explanations, we show the proportions of “similar to ground-truth explanations" in the ECQA dataset, as depicted in Figure 4. We leverage GPT evaluator and well-trained NLI evaluators to determine whether the derived explanations from FaithLM and ground-truth explanations are within similar content or not. The results show that FaithLM’s derived explanations are more consistent with the ground-truth NL explanations, which is indicated by a larger proportion of “similar to ground-truth explanations" generated by FaithLM than baseline methods.
Efficacy of Explanation Trigger Prompts.
We first show prompt optimization curves of explanation trigger prompts on three different datasets, and then leverage the optimal explanation trigger prompts to generate explanations via FaithLM. In the experiments, we randomly select 15 instances from the training dataset in each optimization round, and compute the average fidelity scores of the newly derived trigger prompts. After the progress is terminated, we evaluate the optimized trigger prompts on the testing set. The optimization step is uniformly established at 50 rounds across different explainer and targeted LLM settings. We evaluate the derived explanation trigger prompts under two distinct experimental settings, as shown below.
Figure 5: The fidelity in different optimization steps of the trigger prompts (Algorithm 2) on the ECQA, TrivaQA, and COPA datasets. The fidelity grows higher as the number of steps increases.
  • Trigger Prompt Optimization Curve. Figure 5 demonstrates the optimization curves of three datasets. We display the explainer as GPT-3.5-Turbo and Claude-2 and the explainer as Vicuna-7B. We observe that the optimization curve exhibits a generally ascending trend as the step progresses, interspersed with multiple waves throughout the optimization procedure. This indicates that FaithLM generates better explanation trigger prompts after the optimization. More results of optimization curves on remaining datasets are provided in Appendix E.
  • Explanation Generation by Optimized Trigger Prompts. We utilize the optimized explanation trigger prompts to generate explanations following Algorithm 1. The results are displayed in Figure 6(a), including the experiments conducted using all three datasets with Claude-2 as the explainer and Vicuna-7B as the targeted LLM. We observe that optimized explanation trigger prompts obtain higher fidelity scores than the initial human-crafted trigger prompt in generating explanations. This trend is consistent across all datasets, regardless of whether the explanations are refined by Algorithm 1. Additionally, the optimized trigger prompts are effective in improving the fidelity of the derived explanations with or without the optimization of FaithLM.
A Case Study of  FaithLM.
The case studies illustrate the evolving trend via FaithLM, including derived NL explanations, explanation trigger prompts, and contrary explanations in Appendix I. These studies demonstrate that the explanations generated by FaithLM are informative and readable, which enables humans to understand the reasons behind the decision-making process of LLMs.
4.4 Transferability of Optimized Trigger Prompt (RQ2)
We assess the transferability of ultimately optimized trigger prompts across different unseen datasets within the same domain, as depicted in Figure 6(b). Specifically, we transfer the optimized trigger prompts from the ECQA to the Social-IQA dataset, and from the COPA to the XCOPA datasets, without any additional optimization. Specifically, the Social-IQA dataset is dedicated to commonsense question-answering (similar to the ECQA dataset), while the XCOPA dataset specializes in causal reasoning (similar to the COPA dataset). We adopt the Vicuna-7B as the targeted LLM, and Claude-2 as the explainer on these transfer tasks. The fidelity of the derived NL explanation on the target dataset is shown in Figure 6(b). The optimized trigger prompts show better explanation efficacy than human-crafted prompts when it is transferred from a dataset from a similar domain. This shows that the optimized trigger prompts generated by FaithLM possess a great property of data transferability.
(a) Robustness analytics of trigger prompts
(b) Transferability results of trigger prompts
Figure 6: Assessment on the adaptation of the optimized explanation trigger prompts. Figure (a) reveals the robustness evaluation, and figure (b) illustrates the results on transferability.
(a) Evaluated by pre-trained NLI classifiers
(b) Evaluated by GPT-3.5
Figure 7: Ablation studies on evaluating contrary explanation. The results show that contrary explanations obtain opposite meanings to the derived explanations.
4.5 Ablation Studies on Contrary Explanation (RQ3)
The quality of contrary explanations $\neg\mathcal{E}_{\text{NL}}$ determines the efficacy of FaithLM. We leverage the powerful LLMs as the LLM agent to generate contrary explanations, requesting the delivery of high-quality opposite-meaning outputs from their original NL explanations. In this section, we evaluate the quality of contrary explanations, aiming to observe the semantic differences between the original NL explanations and their contrary explanations. To examine the quality, we employ the GPT classifier and two well-trained NLI classifiers, Roberta-Large and XLNet-Large [31]. We leverage each classifier to distinguish whether the relationship between the “original NL explanations" and “contrary explanations" belong to the category of “similar meaning (entailment)," “dissimilar meaning (contradiction)," or “non-relevant (neutral)." We follow the evaluation settings from [23] on GPT-classifier with evaluation prompt provided in Table 7 of Appendix F.
The results are shown in Figure 6 under the randomly sampled 100 instances from the ECQA and COPA datasets. We observe that the two NLI classifiers achieve up to 86% and 82% in the “dissimilar meanings"" category on the ECQA and COPA datasets, respectively. The empirical results of the GPT-classifier demonstrate that the derived explanations from SelfExp obtain more non-faithful information than FaithLM, which may increase the risk for the LLM agent of Fidelity Evaluator in generating non-relevant information as the contrary explanations. In addition, we provide case studies in Appendix I to demonstrate the informativeness and readability of contrary explanations.
5 Conclusion
In this paper, we introduce FaithLM to explain the decision-making process of LLMs. Specifically, FaithLM employs a fidelity enhancement strategy to progressively refine the fidelity of derived explanations and explanation trigger prompts. Experimental results on three datasets demonstrate the effectiveness of FaithLM in terms of the high fidelity of derived explanations, and better alignment with the ground-truth explanations of the datasets. As for future work, we plan to expand FaithLM to investigate high-stack applications, such as the healthcare domain, where transparency concerns are increasingly prominent due to the widespread adoption of black-box LLMs.
Appendix
Appendix A Details about Datasets
The experiments are conducted on the three NLU datasets. The details of the datasets are provided as follows:
  • ECQA [2]. ECQA is an extension of the CQA dataset [37]. Specifically, based on the CQA dataset, it annotates the positive or negative properties and golden explanations for the QA pairs. Due to API cost budgets, we evaluate our framework on the first 500 instances in the ECQA dataset.
  • TriviaQA LongBench [19]. TriviaQA LongBench (TriviaQA-Long) is a reading comprehension dataset. It includes 300 question-answer-evidence triples sourced from the Longbench [5] dataset [^1] https://huggingface.co/datasets/THUDM/LongBench/. This dataset features question-answer pairs crafted by trivia enthusiasts, accompanied by independently sourced evidence documents, providing supervision for answering these questions.
  • Balanced COPA [35, 20]. The Balanced COPA (COPA) dataset is a collection of 500 questions for commonsense causal reasoning. Each question consists of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
Appendix B Related Work
B.1 Post-hoc Explanation
Post-hoc explanation techniques have undergone significant development and discussion, driven by the widespread adoption of black-box ML models across various data modalities. A multitude of post-hoc algorithms has been introduced from two aspects: local and global explanations [30, 13]. Explanations aim to explain the reasoning behind an individual model for each input instance, while global explanations aim to uncover the overall functioning of a complex model [12]. Considering various purposes of explanation, the explanation techniques mainly showcase the explanation from two perspectives, including feature attributions and counterfactual examples. Feature attribution aims to provide users with important scores for each feature’s impact on model predictions, while counterfactual examples aim to offer alternative instances that explicitly assist users in grasping the model’s decision-making process. In recent years, with the growing proficiency and wide usage of black-box LLMs, especially closed-source LLMs service, post-hoc explanations have become increasingly prominent and have garnered significant attention in NLP research due to the inaccessibility of LLMs’ model weights and structure [46].
B.2 Explainability of LLMs
The majority of explanation efforts in LLM research have centered on delivering explanations. One group of studies calculates importance scores for specific tokens [24, 17], another line of progress generates NL explanations by leveraging the pre-trained LLMs with internal model knowledge sources [21, 9, 28], the other group of work leverages LLMs themselves to generate chain-of-thought (CoT) reasoning [22, 34, 8, 8] as the self-explanations through the one feed-forward inference process. Furthermore, some studies aim to yield counterfactual explanations by pre-trained LLMs to assist users in better understanding the decision-making process from LLMs [7, 8]. Although NL explanations offer fantastic human-understandable insights than token-wise explanations, the explanations can lose their fidelity via one feed-forward inference process of pre-trained LLMs. Unreliability and non-fidelity of NL explanations are still a concern [45, 40]. Given our primary aim of producing faithful explanations, our efforts are concentrated on generating NL explanations to improve the likelihood of accurately representing the decision-making process of LLMs.
B.3 LLMs as Optimizers
LLMs as optimizers is a novel paradigm, describing optimization problems in natural language and utilizing the reasoning capabilities of LLMs for optimizing [44]. Depicting optimization problems in natural language enables the optimization of diverse tasks without defining formal specifications, such as prompt optimization [44, 10, 16], agent learning [36], and model labeling [38]. Based on this optimization paradigm, our work introduces a generative explanation framework with a novel estimation method of sentence-level fidelity.
Appendix C Hyper-parameter Settings of FaithLM
The hyper-parameters of FaithLM are given in Table 2. The configuration for explainers is consistent across Claude-2 and GPT-3.5-Turbo, provided that the parameters are adjustable. Likewise, the settings for predictors are uniform, with Phi-2 and Viucua-13B receiving identical hyperparameter configurations during the experiments conducted in this study.
Dataset ECQA TriviaQA-Long COPA
Fidelity-enhanced Optimization Optimization Steps 20 20 20
Temperature of Predictor LLMs 0.7 0.5 0.7
Temperature of Explainer LLMs 0.9 0.9 0.9
Top-P of Explainer LLMs 0.9 0.9 0.9
Trigger-oriented Optimization Optimization Steps 50 100 100
Sampled Instances 30 30 30
Temperature of Predictor LLMs 0.7 0.5 0.7
Temperature of Explainer LLMs 0.9 0.9 0.9
Top-P of Explainer LLMs 0.9 0.9 0.9
Table 2: Hyper-parameters and optimization settings in FaithLM.
Appendix D Computation Infrastructure and Costs
D.1 Computation Infrastructure
For a fair comparison of testing algorithmic throughput, the experiments are conducted based on the following physical computing infrastructure in Table 3.
Device Attribute Value
Computing infrastructure GPU
GPU model Nvidia-A40
GPU number 1
GPU Memory 46068 MB
Table 3: Computing infrastructure for the experiments.
D.2 Computation Costs
The computational costs associated with xLLM primarily differ from the inference costs of local LLMs and the expenses related to API-accessed LLMs. The computational costs depend on the parameter scale and variants of LLMs used in the xLLM framework, shown in Table 4 and 5.
ECQA TrivaQA COPA
Execution Time (Sec.) $\sim$3 $\sim$5 $\sim$3
Execution Cost (\$) $\sim$0.01 $\sim$0.04 $\sim$0.01
Table 4: Computing costs of FaithLM with GPT-3.5 on each dataset.
bs=32 bs=64 bs=96
Execution Time (Sec.) $\sim$3 $\sim$5 $\sim$3
Memory Cost (GB) $\sim$28GB $\sim$43GB $\sim$59GB
Table 5: Computing costs of FaithLM with Vicuna-7B under different batch size (bs).
Appendix E Additional Experimental results of FaithLM
E.1 Optimization Procedure of derived explanations
We demonstrate more evaluation results on derived explanations from FaithLM. The outcomes depicted in Figure 8 reveal that FaithLM attains notably higher fidelity scores across all three datasets following 20 steps of optimization. Additionally, Figure 8 illustrates the evolution of the optimization process during the generation of explanations.
Figure 8: The fidelity evaluation of derived explanations from FaithLM under different settings of predictors and explainers.
E.2 Robustness Analytics of Configurations in FaithLM
In this section, the robustness test of the explainer LLMs is conducted under the analytics of hyper-parameters that are highly dependent on the outputs of LLMs. We focus on two different hyper-parameters: Temperature and Top-p. The experiments are conducted under the explainer GPT-3.5-Turbo and the predictor Vicuna-7B. We evaluate the following temperatures and top-p of the explainer LLM in the range of {0.3, 0.6, 0.9}. The results are shown in Figure 9. We observe that the explainer LLMs perform inferior when the temperature and Top-p are low, reflecting that the lower exploration of explainer LLM may degrade the optimization ability in explanation generation. The explainer LLMs are encouraged to obtain the temperatures and top-p around 0.9. The small values of the temperatures and top-p may lead to low flexibility in updating new explanations. In contrast, large temperatures and top-p may impact explainer LLMs disobeying the given optimization trajectory. Thus, in the main experiments, all reported performances are under the settings of temperature 0.9 and top-p 0.9, achieving the best performance for generating explanations.
Figure 9: Robustness Analytics of FaithLM on different LLM settings, which are Temperature (left) and Top-p strategies (right).
E.3 Additional Optimization Curve of Explanation Trigger Prompt
We demonstrate more evaluation results on the optimization curve of explanation trigger prompts of FaithLM. The optimization curve shown in Figure 10 generally displays an upward trend with the progression of steps, interspersed with several fluctuations throughout the optimization process. This suggests that FaithLM can successfully generate improved explanation trigger prompts after optimization.
Figure 10: The optimization curve of explanation trigger prompts on ECQA (left), TriviaQA-Long (middle), and COPA dataset (right).
E.4 Additional Experiments on Diverse Domains of Dataset
We further have conducted additional experiments on one new MedMCQA dataset [32] in the healthcare domain. We evaluate the FaithLM using a fidelity assessment under Natural Language Explanation Generation settings. All experimental configurations follow the settings in Section 4.2. The experimental results are shown in the table below. We observe that FaithLM outperforms the baseline method, which is consistent with the experimental results across other domain datasets that were evaluated in our work.
SelfExp Self-consistency FaithLM
Fidelity 0.6956 0.4715 0.9565
Table 6: Additional experimental results on MedMCQA dataset.
E.5 Additional Experiments on Truthfulness Evaluation
We further lunch additional experiments on all evaluators with different evaluation settings to measure the relevance between derived explanations and ground-truth explanations. The evaluators assesses their relevance using a GPT-Score in GPT evaluator on a scale of one to five. If the two provided explanations are classified under the “similar content" category or receive a GPT-Score close to five, this indicates that the derived explanations are highly similar to ground-truth explanations.
Figure 11: Truthfulness evaluation with ground-truth explanation under GPT-score settings.
Appendix F Details of Evaluation Prompt Usage
We provide a listing of the evaluation prompts in Table 7 utilized in assessing the performance of FaithLM. The first row reveals the evaluation prompt on comparing the derived explanation with the ground-truth (GT) explanation in the ECQA dataset in Section 4.3; and the second row demonstrates the evaluation prompt on activating the GPT classifier and the GPT scorer for assessing contrary explanations in Section 4.5.
Evaluation Task Evaluation Prompts
Ground-truth Explanation Given a user instruction and two AI assistant responses, your job is to classify whether the relation of two responses in S1 and S2 belongs to G-1, G-2, or G-3. The meaning of class is as follows: (G-1) relevant contents, (G-2) irrelevant contents, or (G-3) irrelevant contents. Judge responses holistically, paying special attention to whether two responses have similar contents. Judge responses with only ONE class label as your final answer. S1:{derived explanation}. S2:{GT-Explanation}. Please ONLY response your in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
GPT classifier for Contrary Explanations Given a user instruction and two AI assistant responses, your job is to classify whether two responses in S1 and S2 belong to G-1, G-2, or G-3. The meaning of class is as follows: (G-1) same semantic meaning, (G-2) opposite semantic meaning, and (G-3) no relation. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. Judge responses with only ONE class label as your final answer. S1:{derived explanation}. S2:{Contrary Explanations}. Please ONLY respond in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
GPT scorer of Contrary Explanations Given a user instruction and two AI assistant responses, your job is to rate from ONE to FIVE to judge whether two responses in S1 and S2 have the same semantic meaning or not. A FIVE score refers to being totally the same, and ONE score refers to being totally the opposite. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. The judge responds with the rates between ONE and FIVE. S1:{derived explanation}. S2:{Contrary Explanations}. Please ONLY respond to the rate value; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
Table 7: Evaluation Prompts given to GPT-3.5-Turbo used in assessing the efficacy of FaithLM.
Appendix G Details of Prompts Usage in FaithLM
We provide a listing of the prompts in Table 7 utilized in FaithLM in different tasks. The first row demonstrates the initial explanation trigger prompt leveraging in both fidelity-enhanced optimization and trigger-oriented optimization. The second row shows the prompt for the LLM agent to generate the contrary explanations.
Conducted Task Evaluation Prompts
Explanation Generation Please provide objective explanations of why the model generates the answers to the given questions based on your thoughts. Explain the reason why the model provides the answer, no matter if it is wrong or correct. Make sure not to answer the questions or provide any suggestions to better answer the questions by yourself. Q:{Question}. A:{Targeted LLM-generated Answer}.
Contrary Explanations Generation Please generate one example of obtaining the opposite meaning from a given sentence. Make sure you output sentences only. Sentences:{derived explanation}.
Table 8: The example of prompts that are given to two explainer LLMs and LLM agent for contrary explanation.
Appendix H Trajectory System Prompts Usage in FaithLM
We present a detailed listing of the trigger-oriented trajectory prompt in Figure 12 and the explanation-oriented prompt in Figure 13, as utilized within the FaithLM framework.
H.1 Trigger-oriented Trajectory Prompt
[…]
Figure 12: A examples of trigger-oriented trajectory prompt. This prompt populates in both LLM explainers, which are Cluade2 and GPT-3.5-Turbo. The output of FaithLM optimized under trigger-oriented trajectory prompt is append after the Response label.
H.2 NL Explanation-oriented Trajectory Prompt
[…]
Figure 13: A examples of explanation-oriented trajectory prompt. This prompt populates in both LLM explainers, which are Cluade2 and GPT-3.5-Turbo. The output of FaithLM optimized under trigger-oriented trajectory prompt is append after the Response label.
Appendix I Case Studies of FaithLM
I.1 Explanations Generation
We demonstrate the case studies of explanation results generated by FaithLM after fidelity-enhanced optimization in the following two Figure 14 and Figure 15. The results show that FaithLM can generate faithful explanations that truly express the decision-making process of targeted LLMs.
Case Study 1 Question: What is the cause of the Promise? Premise: The movie tickets sold out. Choices: [choice] It was opening day for the movie. [choice] The movie received poor reviews. Ground Truth: The movie received poor reviews. LLM Answer: It was opening day for the movie. (wrong) derived explanation: The model likely selected that answer because opening day tends to have very high demand for movies, often resulting in sold out showings. While there could be other reasons for the sell out, high opening day demand is a reasonable explanation given the limited context provided. Case Study 2 Question: What is the effect of the Promise? Premise: I drilled a hole in the wall. Choices: [choice] A mouse crawled out of the hole. [choice] Dust blew out of the hole. Ground Truth: Dust blew out of the hole. LLM Answer: Dust blew out of the hole. (correct) derived explanation: Dust exiting the hole in the wall is a reasonable effect to expect when drilling into a wall. Without further context, there is no reason to assume a mouse coming out of the hole is likely based solely on the premise that a hole was drilled. The model selected the choice directly supported by the information provided in the premise.
Figure 14: Case Study: derived explanations (Dataset: COPA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)
Case Study 1 Question: Where would you get a dog if you do not have one? Choices: [choice] pet store [choice] outside [choice] neighbor’s house [choice] park [choice] rug Ground Truth: store LLM Answer: store (correct) Derived explanation: The model suggested "pet store" as the place to get a dog if you don’t already have one because pet stores are a common retail business where people can purchase dogs as pets. Case Study 2 Question: Where would you put coins if you want to bring them with you? Choices: [choice] medium [choice] water fountain [choice] wishing well [choice] desk [choice] purse Ground Truth: purse LLM Answer: desk (wrong) Derived explanation: The model may have predicted "desk" because desks often have drawers or containers where coins could be stored temporarily. However, a purse is the most practical option for carrying coins when leaving.
Figure 15: Case Study: Derived explanations (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Phi)
I.2 Explanation Trigger Prompts
The demonstrations in the explanation trigger prompts generated by FaithLM in Figure 16. The results show that FaithLM can generate explanation trigger prompts that lead explainer LLMs to generate explanations and obtain higher fidelity.
[…]
Figure 16: Case Study: Generation of Explanation Trigger Prompts
I.3 Contrary Explanation
We demonstrate results on contrary explanations leveraged in Fidelity Evaluator. The results in Figure 17 show that the contrary explanation obtains the opposite meaning of the given explanation.
Case Study 1 Question: The child wanted to test electricity, so what did he feel as a result? Choices: [choice] cause fire [choice] shock [choice] short fuse [choice] kill if used improperly [choice] produce heat Ground Truth: shock LLM Answer: cause fire (wrong) Given Explanation: The model likely incorrectly associated testing electricity with igniting a fire, rather than understanding that direct contact can cause an electric shock. It failed to comprehend the potential outcomes of unsafe electrical contact. Contrary Explanation: The model correctly predicted that testing electricity would not cause a fire. It likely fully understood that direct contact with electricity can cause a shock rather than ignite a fire. Case Study 2 Question: Where would you put coins if you want to bring them with you? Choices: [choice] medium [choice] water fountain [choice] wishing well [choice] desk [choice] purse Ground Truth: purse LLM Answer: purse (correct) Given Explanation: A purse is a personal item that people often carry with them when going places. It has compartments to store small items like coins, so putting coins in your purse allows you to easily bring them along wherever you go. Contrary Explanation: The purse is not a good place to put coins if you don’t want to bring them with you, because purses are designed for other items, not coins.
Figure 17: Case Study: Contrary Explanation (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)
I.4 Limitations of Learning FaithLM
One significant limitation of FaithLM associated with the carbon emissions during the experiments. The extensive computational power required for optimizing FaithLM leads to considerable energy consumption, which, in turn, results in a significant carbon footprint. As the demand for more sophisticated LLMs continues to grow, so does their environmental impact. This limitation underscores the urgent need to explore and adopt more sustainable practices and technologies in the development and learning FaithLM with fewer optimization steps to mitigate their ecological footprint.

### v4

@@
 Abstract
-Large Language Models (LLMs) have become proficient in addressing complex tasks by leveraging their extensive internal knowledge and reasoning capabilities.
-However, the black-box nature of these models complicates the task of explaining their decision-making processes.
-While recent advancements demonstrate the potential of leveraging LLMs to self-explain their predictions through natural language (NL) explanations, their explanations may not accurately reflect the LLMs’ decision-making process due to a lack of fidelity optimization on the derived explanations.
-Measuring the fidelity of NL explanations is a challenging issue, as it is difficult to manipulate the input context to mask the semantics of these explanations.
-To this end, we introduce FaithLM for explaining the decision of LLMs with NL explanations.
-Specifically, FaithLM designs a method for evaluating the fidelity of NL explanations by incorporating the contrary explanations to the query process.
-Moreover, FaithLM conducts an iterative process to improve the fidelity of derived explanations.
-Experiment results on three datasets from multiple domains demonstrate that FaithLM can significantly improve the fidelity of derived explanations, which also provides a better alignment with the ground-truth explanations.
+Large language models (LLMs) increasingly produce natural language explanations, yet these explanations often lack faithfulness, and they do not reliably reflect the evidence the model uses to decide.
+We introduce FaithLM, a model-agnostic framework that evaluates and improves the faithfulness of LLM explanations without token masking or task-specific heuristics.
+FaithLM formalizes explanation faithfulness as an intervention property: a faithful explanation should yield a prediction shift when its content is contradicted.
+Theoretical analysis shows that the resulting contrary-hint score is a sound and discriminative estimator of faithfulness.
+Building on this principle, FaithLM iteratively refines both the elicitation prompt and the explanation to maximize the measured score.
+Experiments on three multi-domain datasets and multiple LLM backbones demonstrate that FaithLM consistently increases faithfulness and produces explanations more aligned with human rationales than strong self-explanation baselines.
+These findings highlight that intervention-based evaluation, coupled with iterative optimization, provides a principled route toward faithful and reliable LLM explanations.
+Our source code is available at https://github.com/ynchuang/xLLM.
+FaithLM: Towards Faithful Explanations for Large Language Models
+Yu-Neng Chuang1thanks: Correspond to Yu-Neng Chuang <ynchuang@rice.edu>, Vladimir Braverman <vova@cs.jhu.edu> and Xia Hu <xia.hu@rice.edu>   Guanchu Wang1   Chia-Yuan Chang2   Ruixiang Tang1 Shaochen Zhong1Fan Yang3Mengnan Du4Xuanting Cai5 Vladimir Braverman56Xia Hu1 Rice University1  Texas A&M University2  Wake Forest University3 New Jersey Institute of Technology4  John Hopkins University5  Google Research6  Meta Platforms, Inc.7
 1 Introduction
-Large language models (LLMs) exhibit remarkable performance in various natural language processing tasks, such as the GPT4 [1], LLaMA [39], and Claude [4].
+Large language models (LLMs) exhibit remarkable performance in various natural language processing tasks, such as the GPT4 Achiam et al.
+(2023), LLaMA Touvron et al.
+(2023), and Claude AnthropicAI (2023).
 However, these language models are commonly regarded as intricate black-box systems.
 The opacity of their internal mechanisms poses a significant challenge when trying to explain their decision-making process.
-The lack of transparency in LLMs, especially in API-accessed LLM services, inferences contradict the practical requirements of stakeholders and are in opposition to regulatory standards in various domains, such as GDPR [15, 14].
+The lack of transparency in LLMs, especially in API-accessed LLM services, inferences contradict the practical requirements of stakeholders and are in opposition to regulatory standards in various domains, such as GDPR Goodman et al.
+(2017); Floridi (2019).
 The imperative arises to develop explainability mechanisms for LLMs, particularly for their use in high-stakes applications such as healthcare.
+In this work, we focus on “LLM explanation", rather than “LLM reasoning" or “LLM self-refinement", to interpret the model prediction behaviors after providing the final responses.
+(More illustrations of their discrepancy are in Section 2.2).
 Numerous studies have attempted to enhance the transparency of decision-making processes in LLMs by providing natural language (NL) explanations.
-Advanced billion-level LLMs, such as GPT4 [1], mainly leverage their reasoning abilities and internal knowledge for complex downstream problems.
-However, this complexity poses challenges to faithfully explain the underlying reasons behind their decisions with natural language sentences.
-In this manner, recent advancements are struggling to generate reliable NL explanations for interpreting LLMs [45].
-Some work attempt to leverage powerful LLMs [27, 9] with auxiliary information to generate NL sentences or chain-of-thought (CoT) reasoning [22, 34, 8] as the explanations to explain the decision-making process of LLMs.
-Although existing work emerged that LLMs may possess the ability to self-explain [26, 42], or follow chain-of-thoughts to explain [25], their explanation-generating process usually overlooks the fidelity, a fundamental metric for evaluating the quality of explanations [12, 41].
-Thus, the derived NL explanations may not faithfully reflect the decision-making process of LLMs [46, 40].
-Measuring the fidelity of NL explanations is a challenging issue, as the NL explanations may provide crucial information beyond the input context.
-This makes it hard to mask the semantics of NL explanations by manipulating the input tokens.
+However, this complexity poses challenges to faithfully explain the underlying explanation behind their decisions with natural language sentences.
+Recent advancements are struggling to generate reliable NL explanations for interpreting LLMs Ye and Durrett (2022).
+Some work attempt to leverage powerful LLMs Majumder et al.
+(2021); Chen et al.
+(2023c, b) with auxiliary information to generate NL sentences or heatmap of input tokens as model explanations.
+Although existing work emerged that LLMs may possess the ability to self-explain Madsen et al.
+(2024), their explanation-generating process usually overlooks the fidelity, a fundamental metric for evaluating the quality of explanations Chuang et al.
+(2023); Wang et al.
+(2023).
+The derived NL explanations may not faithfully reflect “why model generate this answers." Zhao et al.
+(2023); Turpin et al.
+(2023) Some work attempts to leverage chain-of-thought (CoT) reasoning steps as the post-hoc model explanation Lyu et al.
+(2023); Radhakrishnan et al.
+(2023).
+However, these reasoning steps are not considered model explanations in the context of post-hoc explanation Tanneru et al.
+(2024), where post-hoc ones particularly focus on providing faithful explanations for a given generated answer.
+These CoT steps are produced without a thorough fidelity check (i.e., one that involves masking out the key factors) to ensure they genuinely influence the final answer.
+These steps are only the intermediate results during the LLM prediction and their fidelity remains unknown.
+A proper fidelity measurement requires masking the critical features or key messages in the explanation and observing the model’s performance afterward Du et al.
+(2019), but neither of them are monitored or adopted before claiming CoT reasoning as model explanations.
+Measuring the fidelity of NL explanations now become a important but challenging issue, as we can monitor and optimization the explanation generation process based on fidelity improvement.
+The ones may provide crucial information beyond the input context, but the faithful information may appear in semantic levels, making it hard to measure by manipulating the tokens for fidelity measurement.
 Question and LLM Answer Faithful NL Explanation Question conditioned with Contrary NL Explanation
@@
 Given a faithful explanation “similar poles pull each other away," with its contrary NL explanation, the LLM changes the answer from No to Yes when introduced contrary NL explanation as an extra condition to LLM.
-This example indicates that the contrary explanation interrupts the LLM’s original prediction process with faithful information, demonstrating that the information in an original explanation is faithful and aligns with LLM’s initial answer.
-To overcome this challenge, we propose a method to measure the fidelity of NL explanations.
-We give an example to convey the motivation in Figure 1.
-Specifically, the fidelity of an explanation can be measured by leveraging its contrary explanation as extra conditions of the input context, and observing the LLM’s output difference compared with its initial output.
-Here, a contrary explanation refers to a statement with opposite semantics to the original explanation.
-By incorporating the contrary explanation to the input context, we can identify an explanation as high fidelity if there is a significant change in the LLM’s output, such as from No to Yes.
-This change indicates that the crucial information present in the original explanation is substituted with the opposite meaning context in the contrary explanation, where the crucial information is essential to the LLM’s decision-making process.
-Based on this observation, we propose to extend the applicability of fidelity to the evaluation on NL explanations.
-This extension follows the integration of contrary explanations to represent the concepts of masking important features in traditional fidelity measurement.
-Building upon this approach for measuring the fidelity of NL explanations, we introduce Faithful LLM Explainers (FaithLM) to generate faithful NL explanations for LLMs.
-Specifically, FaithLM adopts LLMs as explainers to generate the NL explanations and explanation trigger prompts, and iteratively optimizes the derived NL explanations and trigger prompts with the goal of fidelity enhancement.
+This example indicates that the contrary hint interrupts the LLM’s original prediction process with faithful information, demonstrating that the information in an original explanation is faithful and aligns with LLM’s initial answer.
+To overcome this challenge, we assess the fidelity of natural language explanations by treating faithfulness as a causal property: if an explanation captures information or hints the model actually uses, then intervening on that information should predictably change the model’s output.
+Concretely, given an explanation, we construct a contrary hint that expresses the opposite semantics and append it to the input.
+We then measure fidelity as the resulting prediction shift relative to the original output.
+A large, directionally consistent change (for example, from No to Yes) indicates that the explanation contained decision-relevant content that the contrary hint displaced.
+This evaluation mirrors recent work Chen et al.
+(2025) that tests whether models use and verbalize externally provided hints, and that reads faithfulness from sensitivity under controlled prompt interventions.
+Our procedure applies the same principle to explanation content, enabling a simple, model-agnostic fidelity measure for free-text rationales
+Building upon this new fidelity measurement, we introduce Faithful LLM Explainers (FaithLM) to generate faithful NL explanations for LLMs.
+Specifically, FaithLM adopts LLMs as explainer to generate the NL explanations and explanation trigger prompts, and iteratively optimizes the derived NL explanations and trigger prompts with the goal of fidelity enhancement.
 During the iterative process, FaithLM computes the fidelity of each derived explanation and optimized prompt based on our proposed fidelity measurement method, and progressively improves their fidelity through in-context learning.
-To demonstrate the effectiveness of FaithLM, we conducted several experiments on four different LLMs under three datasets from different domains.
-FaithLM achieved significantly higher fidelity performance in generating NL explanations and more closely matched the golden explanations compared with state-of-the-art baseline methods.
-In summary, our contributions can be summarized as follows:
-  • Fidelity of NL Explanations: We propose to measure the fidelity of NL explanation by incorporating its contrary explanation into the query and observing LLM output difference.
-  • Faithful LLM Explainers: We introduce FaithLM for improving the fidelity of NL explanations, aiming at faithfully explaining the decision-making process of LLMs.
-  • Fidelity and Truthfulness: Experimental results on three datasets show that FaithLM can effectively improve the fidelity of NL explanations, which also shows a better alignment with the ground-truth explanations than state-of-the-art baseline methods.
+We conducted the experiments on four different LLMs under three datasets.
+FaithLM achieves significantly higher fidelity in generating NL explanations and more closely matched the golden explanations compared with state-of-the-art baseline methods.
+Our contributions can be summarized as follows:
+  • Intervention-based fidelity. We define fidelity as prediction sensitivity to a contrary hint that semantically opposes the explanation.
+  • Faithful LLM Explainers. Our method generates contrary hints, computes fidelity from outputs, and iteratively refines the prompt and explanation to increase faithfulness.
+  • Empirical gains. Experimental results show that FaithLM raises measured fidelity and improves alignment with human explanation over baselines.
 2 Preliminaries
@@
 However, the directly generated $\mathcal{E}_{\text{NL}}$ under single-forward passing may not be faithful and accurate, degrading the user’s trust in the prediction made by the targeted LLM.
-To this end, our work focuses on benefiting the explainer $\textsl{g}(\cdot)$ to generate more faithful NL explanations regarding the decision of $f(\cdot)$ in a post-hoc manner, where $f(\cdot)$ can be either closed-source or open-source LLMs.
-2.2 Post-hoc Explanations of LLMs
-Recent research on LLM explanations has primarily focused on post-hoc approaches due to the inaccessibility of LLM weights and architectures.
-One group of studies calculates importance scores for specific tokens [24, 17], and the other line of work generates natural language sentences [21, 7, 9, 28] and CoT reasoning [25, 8] as explanations by leveraging LLMs.
-Despite the advanced capability of powerful LLMs to generate NL explanations, issues with unreliability and non-fidelity persist.
-While studies like Chen et al.
-[6] utilize internal model weights aiming for reliable explanations generation, accessing these weights is challenging with closed-source LLMs.
-Considering the non-accessibility of LLMs, such as closed-source LLMs, our goal is to faithfully generate NL explanations instead of exploring the internal mechanisms or neurons of LLMs to generate explanations.
-2.3 Limitations of Traditional Fidelity on Natural Language Explanations
+The consistency between $f(\cdot)$ and $\textsl{g}(\cdot)$ is ensured through an iterative optimization process monitored by fidelity scores.
+To this end, the explainer $\textsl{g}(\cdot)$ to generate more faithful NL explanations regarding the decision of $f(\cdot)$ in a post-hoc manner, where $f(\cdot)$ can be either closed-source or open-source LLMs.
+FaithLM targets realistic post-hoc interpretability for API-only or black-box LLMs, so we separate the target model $f(\cdot)$ that answers from the explainer $g(\cdot)$ that proposes $E_{NL}$ (and $\lnot E_{NL}$).
+2.2 Difference between LLM Explanation and Chain-of-thoughts
+Due to the limited accessibility of LLM APIs, recent research on LLM explanations has largely relied on post-hoc explanation approaches (Chen et al., 2023c).
+However, some studies conflate ’LLM reasoning’ and “LLM self-refinement’ with ’LLM explanations’ when discussing these post hoc LLM explanations, even though these three terms are not identical and with different goals.
+We illustrate the difference as follows.
+LLM reasoning and Chain-of-thoughts
+refers to the internal process the model undergoes when it encounters a query or instruction, such as weighting probabilities and generating words step by step plus verification, with the goal of improving performance on reasoning tasks.
+Some advantages rely on providing chain-of-thought (CoT) reasoning (Lanham et al., 2023; Radhakrishnan et al., 2023; Chen et al., 2023b; Wang et al., 2022) to present the hidden inference steps that the model goes through.
+These studies show that CoT manuvinakurike2025thoughts can improve reasoning performance, but does not necessarily provide an explanation or even count as explanations of how or why an LLM arrives at its answers (Tanneru et al., 2024), where “good fidelity” in the series of work is typically defined by the alignment between the content of CoT and the final answer (Lyu et al., 2023; Radhakrishnan et al., 2023).
+Another line of work leverages self-refinement techniques (Lightman et al., 2023; Madaan et al., 2024; Tian et al., 2024), which employ self-reasoning or knowledge supervision as feedback, to iteratively enhance reasoning performance.
+Although these advancements introduce robust self-feedback loops that effectively boost reasoning accuracy, the “feedback” during optimization is neither necessarily faithful nor equivalent to LLM explanations (Tanneru et al., 2024).
+Notably, this feedback may be wrong yet still guide LLMs toward a correct reasoning direction.
+Due to its non-stationary nature, it yields non-faithful outputs when treated as an LLM explanation, which is also very distinct from the goal of “LLM explanation" tasks.
+LLM Explanations.
+Unlike LLM reasoning and self-refinement, LLM explanation focuses on clarifying why the model provides a particular answer after generating its final decision (Siegel et al., 2024).
+The concept of fidelity in an LLM explanation (Du et al., 2019; Zhao et al., 2023), which differs from LLM reasoning and self-refinement, refers to whether the model’s prediction would change if the key knowledge provided explanation were removed.
+If removing the knowledge causes a drastic change in the model’s prediction, we can conclude that the derived explanation is faithful to the LLMs prediction (i.e., the actual reasons that results the predictions).
+In this work, we focus on LLM explanation, rather than LLM reasoning or self-refinement.
+2.3 Limitations of Traditional Fidelity Measurement on NL Explanations
 The fidelity metric measures the fidelity of the given explanation, which is broadly applicable when ground-truth explanations are unavailable.
-In the NLP scenario, fidelity has been used to evaluate the heatmap-formatted explanations [24, 17], where the heatmap one highlights the important tokens of the input.
+In the NLP scenario, fidelity has been used to evaluate the heatmap-formatted explanations Lopardo et al.
+(2023); Huang et al.
+(2023), where the heatmap one highlights the important tokens of the input.
 Specifically, fidelity evaluates the explanation by removing the important tokens from the input $\boldsymbol{X}$ and checking the prediction difference of the targeted LLM.
-Following the definition of fidelity [29].
+Following the definition of fidelity Miró-Nicolau et al.
+(2024).
 Given a sequence of tokens $\boldsymbol{I}=\{t_{1},\cdots,t_{M}\}\subseteq\mathcal{E}_{\text{NL}}$, which is identified as an important component of explanation to the prediction of a targeted LLM $\boldsymbol{Y}=f(\boldsymbol{X})$.
-Following the traditional fidelity definition, the fidelity can be estimated as follows:
+Following the traditional fidelity definition, the fidelity can be estimated as:
 $\mathrm{Fidelity}=f(\boldsymbol{X})-f(\boldsymbol{X}\setminus\boldsymbol{I}),$
-where "$\boldsymbol{X}\setminus\boldsymbol{I}$" denotes the token removal from $\boldsymbol{X}$ in $\boldsymbol{I}$.
+where $\boldsymbol{X}\setminus\boldsymbol{I}$" denotes token removal from $\boldsymbol{X}$ in $\boldsymbol{I}$.
 If important component $\boldsymbol{T}$ achieves higher fidelity, this demonstrates that $\mathcal{E}_{\text{NL}}$ comprises the crucial tokens that significantly influence the predictions of the targeted LLMs.
-However, it is challenging to evaluate the fidelity of NL explanations throughout the fidelity defined above, as the critical components in NL explanations may not contain in the input context $\boldsymbol{X}$.
-For instance, as the example shown in Figure 1, the critical component “similar poles push each other away" in $\mathcal{E}_{\text{NL}}$ does not explicitly appear in the input context.
-Thus, we cannot simply remove critical components from the question following the traditional definition.
-Therefore, in this paper, we propose a solution to systematically address this obstacle by removing the critical components from the semantic level instead of the token level.
+However, it is challenging to evaluate the fidelity of NL explanations throughout the fidelity defined above, as the critical components in NL explanations may not contain in the input context $\boldsymbol{X}$.Some work Lanham et al.
+(2023) attempts to measure fidelity by modifying the output chain-of-thought (CoT) reasoning to overcome this challenge.
+However, altering only the output does not guarantee changes in the model’s pre-filling probability and may therefore meet self-consistency, but rather than the definition of fidelity.
+Thus, we cannot simply remove or modify critical components from the question following the traditional definition.
+Unlike the previous approaches, we propose a solution to systematically address this obstacle by removing the critical components from the semantic level instead of the token level.
 3 FaithLM: The Explainer LLM Framework
-In this section, we systematically introduce the generative explanation framework, FaithLM, which derives faithful explanations in natural language.
-The derived explanations are expected to accurately reflect the predictive decision-making process of targeted LLMs with high fidelity.
-Specifically, we first introduce the Fidelity Evaluator in Section 3.1 to quantitatively assess the fidelity of the NL explanations.
-Subsequently, FaithLM iteratively optimizes the fidelity of the derived explanations and searches for improved explanation trigger prompts using the proposed Fidelity Evaluator.
-The optimization frameworks for these processes are detailed in Sections 3.2 and 3.3.
-3.1 Fidelity Evaluator for Natural Language Explanations
-Due to the limitation of the fidelity metric described in Section 2.3, we introduce the Fidelity Evaluator to directly assess the fidelity of NL explanations, as shown in Figure 1.
-Specifically, we first extend the original fidelity definition to evaluate the fidelity of NL explanations $\mathcal{E}_{\text{NL}}$.
-Finally, we introduce the fidelity evaluator that follows the proposed fidelity measurement criteria.
-Fidelity of NL Explanations.
-Figure 1: The framework of Fidelity Evaluator. The evaluator calculates the fidelity scores of the derived explanations based on its contrary explanations.
-To assess the fidelity of NL explanations, we extend the traditional fidelity definition to equip it with special constraints regarding NL explanations.
-To address this challenge, we propose formulating the fidelity of NL explanation as the prediction difference caused by the “contrary explanation".
-Specifically, the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ is defined as a statement obtained opposite semantics to the given NL explanation $\mathcal{E}_{\text{NL}}$.
-For instance, if the explanation is "similar poles push each other away," then the contrary explanation would be "similar poles pull each other closer." To estimate the fidelity of $\mathcal{E}_{\text{NL}}$, we use the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ as an extra given condition to the input queries, forming conditional LLM inferences $f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$.
-This operation results in different prediction results for targeted LLMs compared to the original predictions, which compels the targeted LLMs to follow contrary information from the explanations of the input queries.
-In this manner, the fidelity $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$ can be estimated by the prediction difference:
-$\mathcal{S}_{\mathcal{E}}:=f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}}).$ (1)
-This formulation aligns with the traditional definition of fidelity by observing the prediction difference of the LLM when considering input with and without the component supplied by the explanations.
-Fidelity Evaluator.
-Unreliable explanations derived from LLMs often present incorrect predictions [45], which means that the relationship between the reliability of the explanations and the correctness of predictions is significantly high.
-Motivated by the observation, we propose a framework of the Fidelity Evaluator as illustrated in Figure 1.
-Specifically, we utilize a powerful LLM agent, such as GPT-3.5, to generate the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ regarding the original $\mathcal{E}_{\text{NL}}$.
-The generation process for $\neg\mathcal{E}_{\text{NL}}$ is guided by the prompt provided in Appendix G.
-Given the contrary explanation as a condition for the input context, the Fidelity Evaluator computes the fidelity scores according to the difference of output logits $f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$.
-Intuitively, if the targeted LLMs’ output probability changes significantly enough to flip the result (i.e., judgements or choices), it implies that the contrary explanations $\neg\mathcal{E}_{\text{NL}}$ consist of the contrasting key messages against $\mathcal{E}_{\text{NL}}$ that can significantly impact the decision-making process of the targeted LLM.
-This suggests that the explanation $\mathcal{E}_{\text{NL}}$ contains key components that significantly support the inference process of the targeted LLM.
-3.2 FaithLM for Fidelity-enhanced Explanation
+In this section, we systematically introduce the generative explanation framework, FaithLM, which derives faithful explanations in natural language format.
+The derived explanations are expected to accurately reflect the predictive decision-making process of targeted LLMs with high fidelity after optimizing under FaithLM.
+Figure 1: The fidelity evaluation with hint. The evaluator calculates the fidelity scores of the derived explanations based on its contrary hints.
+(a) Fidelity-enhanced Explanation
+(b) Trigger Prompt Optimization
+Figure 2: An overview of FaithLM framework for two differenent optimization objectives.
+The blue dotted line reveals the trajectory to optimize the NL explanation (Section 3.2), and the red dotted line indicates the trajectory of the explanation trigger prompt optimization (Section 3.3).
+“Traj.
+Prompt" denotes the trajectory system prompt shown in Appendix J.
+3.1 Fidelity via Contrary Hint Interventions
+To reduce penalization from irrelevant tokens, FaithLM computes the probability gap $\Delta p(Y|X)$ only over tokens with non-zero attribution in $f(\cdot)$’s saliency map, thereby isolating fidelity shifts caused by semantically relevant regions.
+Contrary Hint Interventions
+We evaluate the fidelity of a natural language explanation $\mathcal{E}_{\text{NL}}$ by treating faithfulness as an intervention property on model inference, in line with recent work that tests faithfulness via controlled changes to the information a model conditions on (Chen et al., 2025; Lanham et al., 2023).
+Let $f(\mathbf{X})$ denote the model prediction on input $\mathbf{X}$.
+We construct a contrary hint $\neg\mathcal{E}_{\text{NL}}$, defined as a statement whose semantics are opposed to $\mathcal{E}_{\text{NL}}$ (for example, “similar poles pull each other closer” as the contrary of “similar poles push each other away”).
+We then condition the model on this contrary hint and read out a fidelity score from the intervention-induced prediction shift:
+$\mathcal{S}_{\mathcal{E}}:=f(\mathbf{X})-f(\mathbf{X}\mid\neg\mathcal{E}_{\text{NL}}).$
+If $\mathcal{E}_{\text{NL}}$ captures information the model actually uses, inserting $\neg\mathcal{E}_{\text{NL}}$ produces a measurable and directionally consistent change in the output distribution, often sufficient to alter the discrete decision.
+This operationalization supports free-text explanations without token masking or task-specific heuristics.
+In practice, our Fidelity Evaluator (i) prompts a strong LLM to produce $\neg\mathcal{E}_{\text{NL}}$ from $\mathcal{E}_{\text{NL}}$ under a fixed instruction template and (ii) estimates $\mathcal{S}_{\mathcal{E}}$ from output logits or calibrated probabilities, using the sensitivity to the contrary-hint intervention as the fidelity signal.
+Relation to prior faithfulness evaluations.
+Our evaluation protocol follows the intervention-based paradigm introduced by prior work (Chen et al., 2025) to test whether verbalized rationales genuinely reflect a model’s internal decision process (Chen et al., 2025).
+That work manipulates the presence and semantics of hints in prompts and measures whether models both use and verbalize those hints, revealing that stated chains of thought are often unfaithful to underlying reasoning.
+We adopt the same causal-testing philosophy but apply it to explanation content rather than reasoning text.
+Specifically, the contrary hint serves as a controlled semantic intervention on the information encoded in the explanation, and the resulting change in model prediction quantifies its fidelity.
+In this formulation, faithfulness is defined as prediction sensitivity: an explanation is faithful if altering its truth value changes the model’s decision.
+Consequently, our contrary-hint score provides a direct, model-agnostic instantiation of the evaluation approach advocated by Chen et al.
+(2025), complementing prior intervention tests on CoT text (Lanham et al., 2023) by operating at the level of meaning rather than surface form.
+Theorem 1 (Latent-Context Intervention Validity for Faithfulness).
+Let $f:\mathcal{X}\!\times\!\mathcal{C}\!\to\!\Delta(\mathcal{Y})$ be a language model mapping an input $X$ and latent context $C$ to a predictive distribution over an output space $\mathcal{Y}$.
+Let $E_{NL}$ denote a natural-language explanation of $f(X;C)$, and let $\lnot E_{NL}$ denote its contrary hint.
+Assume that $E_{NL}$ asserts a proposition about a semantic factor $S_{E}=s(X,C)$, where $s(\cdot)$ extracts the decision-relevant concept, which is latent or retrieved, that the explanation verbalizes.
+Conditioning on $\lnot E_{NL}$ is equivalent to intervening on this factor while holding $(X,C)$ fixed, i.e., $f(X;C\mid\lnot E_{NL})=f(X;C\mid do(S_{E}\!\leftarrow\!\bar{s}))$ for some contradictory value $\bar{s}$, and predictions are invariant to any irrelevant text $R$, so $f(X;C)=f(X\cup R;C)$.
+Defining $S_{E}(X;C)=D(f(X;C),f(X;C\mid\lnot E_{NL}))$, where $D$ is any strictly proper divergence, we have
+$S_{E}(X;C)=0$ $\iff E_{NL}\text{ is non-faithful for }f(X;C),$
+$S_{E}(X;C)>0$ $\iff E_{NL}\text{ is faithful for }f(X;C).$
+Hence, the contrary-hint score $S_{E}$ constitutes a valid empirical estimator of faithfulness when the decision-relevant content is not contained in the observed input $X$ but arises from latent or retrieved context.
+Intuition of the theoretical foundation.
+Theorem 1 formalizes this causal perspective: if the explanation encodes information that lies on a causal path to the model’s output, then intervening with a contradictory hint necessarily shifts the predictive distribution, yielding a positive contrary-hint score.
+Conversely, if the explanation is merely correlational or decorative, the intervention leaves the model unchanged and the score remains zero.
+Thus, the contrary-hint fidelity measure $S_{E}$ causal faithfulness by directly testing whether the model’s output is functionally dependent on the semantics expressed in its own explanation.
+More discussions and proof are in Appendix A and Corollary 1.
+3.2 FaithLM on Fidelity-enhanced Explanation
 In this section, we introduce an iterative framework designed to progressively enhance the fidelity of NL explanations.
@@
 We aim to explain the response $\boldsymbol{Y}$ produced by the targeted LLM $f(\cdot)$ in response to the given input queries $\boldsymbol{X}$ with NL explanations $\mathcal{E}_{\text{NL}}$ following the goal of fidelity enhancement.
-In the first round of enhancement, the LLM explainer generates NL explanations $\mathcal{E}_{\text{NL}}$ following human-crafted explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ provided in Appendix G.
+In the first round of enhancement, the LLM explainer generates NL explanations $\mathcal{E}_{\text{NL}}$ following a given human-crafted explanation trigger prompt $\mathcal{P}_{\mathcal{E}}$ provided in Appendix I.
 The explanations are then generated by the explainer $\textsl{g}(\mathcal{P}_{\mathcal{E}}|~\boldsymbol{X},\boldsymbol{Y})$.
@@
 Following this trajectory, the LLM explainer generates new explanations with the goal of achieving higher fidelity scores in subsequent iterations.
-This fidelity enhancement process is guided by the system prompts detailed in Figure 13 of Appendix H.
+This process is guided by the system prompts detailed in Figure 14.
 The trajectory $\mathcal{T}$ is continuously updated by incorporating each newly derived explanation with its assessed fidelity score until the convergence.
 Regardless of any given explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$, FaithLM can all systematically guide the generation of NL explanations, progressively improving fidelity scores by following the reference path established in the trajectory.
-(a) Fidelity-enhanced Explanation
-(b) Trigger Prompt Optimization
-Figure 2: An overview of FaithLM framework for two differenent optimization objectives.
-The blue dotted line reveals the trajectory to optimize the NL explanation (Section 3.2), and the red dotted line indicates the trajectory of the explanation trigger prompt optimization (Section 3.3).
-“Traj.
-Prompt" denotes the trajectory system prompt shown in Section H.
-Algorithm 1 FaithLM for fidelity-enhanced explanation
-Input: Input $\boldsymbol{X}$, output $\boldsymbol{Y}$, targeted LLMs $f(\cdot)$, human-crafted prompt $\mathcal{P}_{\mathcal{E}}$, and LLM explainer $\textsl{g}(\cdot)$.
-Output: NL explanation $\mathcal{E}_{\text{NL}}$.
-1:  $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}~{}|~{}\boldsymbol{X},\boldsymbol{Y})$
-2:  $\mathcal{T}=\varnothing$
-3:  while  steps not end and decision not flips  do
-4:     Estimate the fidelity score $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$
-5:     Append $\mathcal{T}\leftarrow~{}\mathcal{T}\cap[\mathcal{E}_{\text{NL}},~{}\mathcal{S}_{\mathcal{E}}]$
-6:     Update $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{T}~{}|~{}\boldsymbol{X},\boldsymbol{Y})$
-7:  end while
 Algorithm of Fidelity-enhanced Explanation.
@@
 The iteration terminates at a predetermined step or ceases earlier as soon as FaithLM observes a flipping performance from the targeted LLM $f(\cdot)$.
-3.3 FaithLM for Trigger Prompt Optimization
+Algorithm 1 Fidelity-enhanced Explanation
+Input: Input $\boldsymbol{X}$, output $\boldsymbol{Y}$, targeted LLMs $f(\cdot)$, human-crafted prompt $\mathcal{P}_{\mathcal{E}}$, and LLM explainer $\textsl{g}(\cdot)$.
+Output: NL explanation $\mathcal{E}_{\text{NL}}$.
+[…]
+[…]
+3:while steps not end and decision not flips do
+4:  Estimate the fidelity score $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$
+5:  Append $\mathcal{T}\leftarrow~\mathcal{T}\cap[\mathcal{E}_{\text{NL}},~\mathcal{S}_{\mathcal{E}}]$
+6:  Update $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{T}~|~\boldsymbol{X},\boldsymbol{Y})$
+7:end while
+3.3 FaithLM on Trigger Prompt Optimization
 Despite the success of enhancing fidelity in Section 3.2, the low quality of the explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ may still hinder the optimization process of receiving a high-fidelity explanation.
 Given that the unknown preference for prompts from LLMs, human-crafted trigger prompts used in Fidelity-enhanced Explanation Optimization might lead to sub-optimal fidelity enhancement in the derived explanations.
-In this section, we hereby propose a new optimization pipeline under FaithLM, aiming to optimize the trigger prompt $\mathcal{P}_{\mathcal{E}}$ for generating explanations with higher fidelity.
+In this section, we hereby propose a new optimization pipeline under FaithLM, aiming to optimize the trigger prompt $\mathcal{P}_{\mathcal{E}}$ for generating NL explanations with higher fidelity scores as the LLM explanations of input each input query.
 Trigger Prompt Optimization.
@@
 Different from the optimization goal in Section 3.2, the trajectory in this task collects the trigger prompts $\mathcal{P}_{\mathcal{E}}$ and their fidelity scores $\mathcal{S}_{\mathcal{P}}$.
-The trajectory is constructed by the system optimization prompts detailed in Figure 12 of Appendix H.
+The trajectory is constructed by the system optimization prompts detailed in Figure 13.
 To estimate the fidelity score for a trigger prompt, FaithLM first adopts the randomly human-crafted trigger prompt to guide the LLM explainers to generate NL explanations, and then utilize the Fidelity Evaluator to assess the fidelity of the derived explanation.
-The final estimated score is averaged by the fidelity score $\mathcal{S}_{\mathcal{E}_{i}}$ of each instance $i$ from the hold-out dataset $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\in\mathcal{D}$.
-Formally, the fidelity score for a trigger prompt $\mathcal{P}_{\mathcal{E}}$ can be estimated as follows:
+The final estimated score is averaged by the fidelity score $\mathcal{S}_{\mathcal{E}_{i}}$ of the hold-out dataset $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\in\mathcal{D}$.
+Formally, the fidelity score for a trigger prompt $\mathcal{P}_{\mathcal{E}}$ is as follows:
 $\mathcal{S}_{\mathcal{P}}=\mathbb{E}_{\mathcal{E}_{i}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})}\big[\mathcal{S}_{\mathcal{E}_{i}}\big],$ (1)
-where $\mathcal{S}_{\mathcal{E}{i}}$ represents the fidelity score of the explanation $\mathcal{E}_{i}$, which is generated by $\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$, as assessed by the Fidelity Evaluator discussed in Section 3.1.
-During the optimization procedure, the trajectory begins from an empty set and starts to incorporate newly derived trigger prompts along with the fidelity scores in each optimization iteration.
-Following this trajectory, the LLM explainer generates a new trigger prompt with the goal of achieving higher fidelity scores of explanations in subsequent iterations.
-After several rounds of iterations, FaithLM ultimately yields an optimal explanation trigger prompt with the highest fidelity score for the LLM explainer to generate a more faithful NL explanation.
+where $\mathcal{S}_{\mathcal{E}{i}}$ represents the fidelity score of the explanation $\mathcal{E}_{i}$, which is generated by $\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$, as assessed by the Fidelity Evaluator.
 Algorithm 2 Trigger Prompt Optimization.
@@
 1: Initialize human-crafted $\mathcal{P}_{\mathcal{E}}$
-[…]
+2: Initialize $\mathcal{T}=\{\varnothing$}
 3:while (Steps Not End) do
@@
 11:end while
+During the optimization, the trajectory begins from an empty set and starts to incorporate newly derived trigger prompts with the fidelity scores in each optimization iteration.
+Following this trajectory, the LLM explainer generates a new trigger prompt with the goal of achieving higher fidelity scores of explanations in subsequent iterations.
+After several rounds of iterations, FaithLM ultimately yields an optimal explanation trigger prompt with the highest fidelity score for the LLM explainer to generate a more faithful NL explanation.
 Algorithm of Trigger Prompt Optimization.
@@
 Through multiple iterations, FaithLM progressively guides the trigger prompt to generate explanations with higher fidelity scores, following the reference path established in the trajectory.
-The iteration process terminates at a predetermined step, and we set 20 in our experiments.
+The iteration process terminates at a predetermined 20 step.
+Figure 3: The fidelity evaluation of explanations on ECQA (left), TriviaQA-Long (middle), and COPA dataset (right).
+The reported scores are the average fidelity on testing instances in each step of fidelity-enhanced optimization.
 4 Experiment
@@
 Datasets.
-We evaluate FaithLM on three datasets with multiple tasks: ECQA [2] dataset on commonsense question-answer task, TrivaQA-Long [5, 19] dataset on reading comprehension task, and COPA [20, 35] dataset on commonsense causal reasoning task.
-More details of datasets are provided in Appendix A.
+We evaluate FaithLM on three datasets with multiple tasks: ECQA Aggarwal et al.
+(2021) dataset on commonsense question-answer task, TrivaQA-Long Bai et al.
+(2023); Joshi et al.
+(2017) dataset on reading comprehension task, and COPA Kavumba et al.
+(2019); Roemmele et al.
+(2011) dataset on commonsense causal reasoning task.
+More details of datasets are provided in Appendix B.
 Baseline Methods.
-We compare FaithLM with two state-of-the-art baseline methods: SelfExp [26] and Self-consistency [42].
+We compare FaithLM with two state-of-the-art baseline methods: SelfExp Madsen et al.
+(2024) and Self-consistency Wang et al.
+(2022).
 The former ones instruct LLMs to generate explanations using prompt engineering under single-forward inference, and the later ones leverage the outputs from the chain-of-thought prompting process as the model explanations.
-4.2 Experiment Settings
-In this part, we introduce the experimental settings for evaluating FaithLM. Two distinct types of explanation tasks and evaluation settings are illustrated as follows.
+Figure 4: Trustfulness evaluation of the NL explanations.
+Higher the proportion of “similar to ground-truth explanation," the more consistent the derived explanations are with the ground-truth NL explanations.
+Figure 5: The fidelity in different optimization steps of the trigger prompts (Algorithm 2) on the ECQA, TrivaQA, and COPA datasets. The fidelity grows higher as the number of steps increases.
+Comparison with Concurrent Faithfulness Notions.
+Simulatability-based evaluations (e.g.,  Madsen et al.
+(2024); Chen et al.
+(2023a)) assess whether a model can reproduce predictions from an explanation, i.e., behavioral imitation.
+In contrast, FaithLM measures causal sensitivity via a controlled intervention on the explanation’s semantics: we compute $S_{E}=f(X)-f(X\,|\,\lnot E_{NL})$ and read off prediction shifts induced by a contrary hint, without token masking or task-specific heuristics.
+This aligns with intervention-driven faithfulness testing advocated by recent work and keeps the target model $f$ unchanged.
+[^1] See our definition and evaluator description in §3.1.
+Empirically, we additionally report a small-scale comparison where we compute (i) simulatability scores and (ii) $S_{E}$ on the same instances; we analyze where the two agree and where $S_{E}$ detects causal use beyond imitability.
+4.2 Experimental Settings
+We evaluate FaithLM through two complementary tasks: (i) fidelity-enhanced explanation generation and (ii) explanation-trigger prompt optimization. Details are presented in Appendix C.
+Fidelity-Enhanced Explanation.
+We measure the alignment between generated explanations and predictions Chen et al.
+(2025).
+For each example, FaithLM produces one NL explanation, and the average fidelity score across instances is reported as the final metric.
+Explanation-Trigger Prompt Optimization.
+This task optimizes the trigger prompt that guides FaithLM to generate more faithful explanations.
+At each iteration, 30 samples are drawn from the training data as a validation set, and the mean fidelity score on these samples is used as the optimization objective.
+Evaluation Metrics.
+We evaluate explanation quality using two metrics: fidelity and truthfulness.
+Fidelity follows the intervention-based protocol of Chen et al.
+(2025), where contrary hints act as controlled semantic interventions and fidelity is the resulting prediction sensitivity.
+Truthfulness measures semantic consistency between generated and ground-truth explanations using GPT-4o, RoBERTa-Large, and XLNet-Large (Nie et al., 2020), following Liu et al.
+(2023).
+Evaluators classify each pair as similar, dissimilar, or non-relevant, and the proportion of “similar” outputs forms the truthfulness score.
+Full prompts appear in Appendix H.
+Implementation Details.
+We use Vicuna-7B (Chiang et al., 2023) and Phi-2 (Javaheripi and Bubeck, 2023) as target models $f(\cdot)$, and GPT-3.5-Turbo and Claude-2 (Anthropic, 2023) as explainers $g(\cdot)$.
+The same models generate contrary hints.
+Results are averaged over three runs with grid search on hyperparameters.
+Both Vicuna-7B and Phi-2 share identical decoding settings.
+Full configurations and infrastructure details are given in Appendices E and F.
+4.3 Explanation Efficacy of FaithLM (RQ1)
+Efficacy of Derived Explanations.
+We assess the efficacy of derived explanations under the fidelity metric.
+FaithLM adopts the trajectory system prompts in Figure 14 of Appendix J.
+The generation of contrary hints is guided by the prompt in Table 8.
+(a) Robustness analytics of trigger prompts
+(b) Transferability results of trigger prompts
+Figure 6: Assessment on the adaptation of the optimized explanation trigger prompts. Figure (a) reveals the robustness evaluation, and figure (b) illustrates the results on transferability.
+(a) Evaluated by GPT-4o Evaluator
+Figure 7: Ablation studies on evaluating contrary hint.
+  • Fidelity Evaluation.
+The results in Figure 3 demonstrate that FaithLM achieves significantly higher fidelity scores across all three datasets compared with two baselines after 20 steps of optimization.
+Moreover, the optimization curve of fidelity demonstrates that 20 rounds of optimization are sufficient to converge.
+A similar phenomenon occurs across different settings of explainers and targeted LLMs.
+Additional results are provided in Appendix G.
+  • Truthfulness Evaluation.
+To evaluate the truthfulness of explanations, we show the proportions of “similar to ground-truth explanations" in the ECQA dataset, as depicted in Figure 4.
+We leverage GPT evaluators and well-trained NLI evaluators to assess whether the given explanations are within similar content to ground-truth explanations.
+The results show that FaithLM’s explanations are more consistent with the ground-truth NL explanations, indicated by a larger proportion of “similar to ground-truth explanations" generated by FaithLM than baseline methods.
+Efficacy of Explanation Trigger Prompts.
+We first show prompt optimization curves on three different datasets, and then leverage the optimal explanation trigger prompts to generate explanations via FaithLM.
+In the experiments, we randomly select 15 instances from the training dataset in each optimization round, and compute the average fidelity scores of the newly derived trigger prompts.
+After the progress is terminated, we evaluate the optimized trigger prompts on the testing set.
+The optimization step is uniformly established at 50 rounds across different explainer and targeted LLMs.
+  • Trigger Prompt Optimization Curve.
+Figure 5 demonstrates the optimization curves of three datasets.
+We display the explainer as GPT-3.5-Turbo and Claude-2 and the explainer as Vicuna-7B.
+We observe that the optimization curve exhibits a generally ascending trend as the step progresses, interspersed with multiple waves throughout the optimization procedure.
+This indicates that FaithLM generates better explanation trigger prompts after the optimization.
+More results of optimization curves on remaining datasets are provided in Appendix G.
+  • Explanation Generation by Optimized Trigger Prompts.
+We utilize the optimized explanation trigger prompts to generate explanations following Algorithm 1.
+The results are displayed in Figure 6(a), including the experiments conducted using all three datasets with Claude-2 as the explainer and Vicuna-7B as the targeted LLM.
+We observe that optimized explanation trigger prompts obtain higher fidelity scores than the initial human-crafted trigger prompt in generating explanations.
+This trend is consistent across all datasets, regardless of whether the explanations are refined by Algorithm 1.
+4.4 Transferability of Trigger Prompt (RQ2)
+We assess the transferability of ultimately optimized trigger prompts across different unseen datasets within the same domain, as depicted in Figure 6(b).
+Specifically, we transfer the optimized trigger prompts from the ECQA to the Social-IQA dataset, and from the COPA to the XCOPA datasets, without any additional optimization.
+Specifically, the Social-IQA dataset is dedicated to commonsense question-answering (similar to the ECQA dataset), while the XCOPA dataset specializes in causal reasoning (similar to the COPA dataset).
+We adopt the Vicuna-7B as the targeted LLM, and Claude-2 as the explainer on these transfer tasks.
+The fidelity of the derived NL explanation on the target dataset is shown in Figure 6(b).
+The optimized trigger prompts show better explanation efficacy than human-crafted prompts when it is transferred in similar domain.
+This shows that the optimized trigger prompts generated by FaithLM possess a great data transferability.
+4.5 Ablation Studies on contrary hint (RQ3)
+The quality of contrary hints $\neg\mathcal{E}_{\text{NL}}$ determines the efficacy of FaithLM.
+We leverage the powerful LLMs as the LLM agent to generate contrary hints, requesting the delivery of high-quality opposite-meaning outputs from their original NL explanations.
+In this section, we evaluate the quality of contrary hints, aiming to observe the semantic differences between the original NL explanations and their contrary hints.
+To examine the quality, we employ one GPT-4o classifier and two well-trained NLI classifiers, Roberta-Large and XLNet-Large Nie et al.
+(2020).
+We leverage each classifier to distinguish whether the relationship between the “original NL explanations" and “contrary hints" belong to the category of “similar meaning (entailment)," “dissimilar meaning (contradiction)," or “non-relevant (neutral)." We follow the evaluation settings from Liu et al.
+(2023) on GPT-classifier with evaluation prompt provided in Table 7.
+The results are shown in Figure 7 and Figure 11 under the randomly sampled 100 instances from the ECQA and COPA datasets.
+We observe that the two NLI classifiers achieve up to 86% and 82% in the “dissimilar meanings"" category on the ECQA and COPA datasets, respectively.
+The results of the GPT-classifier demonstrate that the derived explanations from SelfExp obtain more non-faithful information than FaithLM, risking the LLM agent of Fidelity Evaluator in generating non-relevant information as the contrary hints.
+Case studies are provided in Appendix K to show the informativeness and readability of contrary hints.
+5 Conclusion
+In this paper, we introduce FaithLM to explain the decision-making process of LLMs, instead of providing reasoning or self-refinement feedback as model explanation.
+Specifically, FaithLM employs a fidelity enhancement strategy to progressively refine the fidelity of derived explanations and explanation trigger prompts.
+FaithLM conducts an iterative process to improve the fidelity of derived explanations.
+Theorem 1 establishes the contrary-hint score as a valid measure of faithfulness.
+Experimental results demonstrate the effectiveness of FaithLM, and better alignment with the ground-truth explanations.
+This suggest that the decision-making process are truly reflected.
+For future work, we plan to extend FaithLM in healthcare, where the needs for transparency is critical given the growing reliance on black-box LLMs.
+6 Limitations
+One significant limitation of FaithLM associated with the carbon emissions during the experiments.
+To generate better fidelity and truthfulness of the derived explanations, FaithLM requires iterating a few rounds of optimization during the explanation derivation.
+This leads to extra computational resources to proceed.
+The extra computational power is thus required for optimizing FaithLM leads to considerable energy consumption, which, in turn, results in a significant carbon footprint.
+As the demand for more sophisticated LLMs continues to grow, so does their environmental impact.
+This limitation underscores the urgent need to explore and adopt more sustainable practices and technologies in the development and learning FaithLM with fewer optimization steps to mitigate their ecological footprint.
+Appendix
+Appendix A Proof of Theorem
+Theorem 1 (Latent-Context Intervention Validity for Faithfulness).
+Let $f:\mathcal{X}\!\times\!\mathcal{C}\!\to\!\Delta(\mathcal{Y})$ be a language model mapping an input $X$ and latent context $C$ to a predictive distribution over an output space $\mathcal{Y}$.
+Let $E_{NL}$ denote a natural-language explanation of $f(X;C)$, and let $\lnot E_{NL}$ denote its contrary hint.
+Assume that $E_{NL}$ asserts a proposition about a semantic factor $S_{E}=s(X,C)$, where $s(\cdot)$ extracts the decision-relevant concept, which is latent or retrieved, that the explanation verbalizes.
+Conditioning on $\lnot E_{NL}$ is equivalent to intervening on this factor while holding $(X,C)$ fixed, i.e., $f(X;C\mid\lnot E_{NL})=f(X;C\mid do(S_{E}\!\leftarrow\!\bar{s}))$ for some contradictory value $\bar{s}$, and predictions are invariant to any irrelevant text $R$, so $f(X;C)=f(X\cup R;C)$.
+Defining $S_{E}(X;C)=D(f(X;C),f(X;C\mid\lnot E_{NL}))$, where $D$ is any strictly proper divergence, we have
+$S_{E}(X;C)=0$ $\iff E_{NL}\text{ is non-faithful for }f(X;C),$
+$S_{E}(X;C)>0$ $\iff E_{NL}\text{ is faithful for }f(X;C).$
+Hence, the contrary-hint score $S_{E}$ constitutes a valid empirical estimator of causal faithfulness even when the decision-relevant content is not contained in the observed input $X$ but arises from latent or retrieved context.
+Proof.
+By the causal definition of faithfulness, an explanation is faithful if and only if altering the truth value of the decision-relevant semantic factor changes the model’s prediction while holding other causes fixed.
+Because $E_{NL}$ asserts a proposition about $S_{E}=s(X,C)$, and conditioning on $\lnot E_{NL}$ implements $do(S_{E}\!\leftarrow\!\bar{s})$ with $(X,C)$ fixed, we have $f(X;C\mid\lnot E_{NL})=f\!\big(X;C\,\big|\,do(S_{E}\!\leftarrow\!\bar{s})\big)$.
+If $E_{NL}$ is non-faithful, then $S_{E}$ has no causal influence on $Y$ under $(X,C)$ and the intervention leaves the predictive distribution unchanged, yielding $S_{E}(X;C)=0$.
+If $E_{NL}$ is faithful, $S_{E}$ lies on a causal path to $Y$ under $(X,C)$ and the intervention changes the predictive distribution; strict propriety of $D$ implies $S_{E}(X;C)>0$.
+Invariance under irrelevant spans follows by the stated stability condition and applies to both terms inside $D$.
+∎
+Corollary 1 (Robustness and Monotonicity).
+Under contextual stability, the contrary-hint score remains invariant to irrelevant input variations, satisfying
+$S_{E}(X;C)=S_{E}(X\cup R;C)\quad\text{for any semantically irrelevant }R.$
+Furthermore, if an iterative procedure produces a sequence of explanations $\{E_{NL}^{(t)}\}$ such that $S_{E}^{(t+1)}(X;C)\geq S_{E}^{(t)}(X;C)$ for all iterations prior to convergence or a decision flip, then the sequence is non-decreasing in causal faithfulness.
+Consequently, any iteration with $S_{E}^{(t)}(X;C)>0$ corresponds to a faithful explanation.
+Appendix B Details about Datasets
+The experiments are conducted on the three NLU datasets. The details of the datasets are provided as follows:
+  • ECQA Aggarwal et al.
+(2021).
+ECQA is an extension of the CQA dataset Talmor et al.
+(2019).
+Specifically, based on the CQA dataset, it annotates the positive or negative properties and golden explanations for the QA pairs.
+Due to API cost budgets, we evaluate our framework on the first 500 instances in the ECQA dataset.
+  • TriviaQA LongBench Joshi et al.
+(2017).
+TriviaQA LongBench (TriviaQA-Long) is a reading comprehension dataset.
+It includes 300 question-answer-evidence triples sourced from the Longbench Bai et al.
+(2023) dataset [^2] https://huggingface.co/datasets/THUDM/LongBench/.
+This dataset features question-answer pairs crafted by trivia enthusiasts, accompanied by independently sourced evidence documents, providing supervision for answering these questions.
+  • Balanced COPA Roemmele et al.
+(2011); Kavumba et al.
+(2019).
+The Balanced COPA (COPA) dataset is a collection of 500 questions for commonsense causal reasoning.
+Each question consists of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
+Appendix C Experiment Settings
+We introduce the experimental settings for evaluating FaithLM. Two distinct types of explanation tasks and evaluation settings are as follows.
 Fidelity-enhanced Explanation
-In this task, our goal is to produce NL explanations that exhibit a higher degree of fidelity.
+In this task, our goal is to produce NL explanations that exhibit a higher fidelity.
 The fidelity is exploited as a metric to evaluate fidelity.
@@
 The quality of the derived NL explanation is evaluated under the fidelity and truthfulness metrics.
-The fidelity follows Section 3.1, which observes the flipping rate of the targeted LLMs by incorporating contrary explanations to the input.
-The evaluation of truthfulness assesses the correlation between the derived NL explanations to the ground-truth explanations Specifically, we leverage GPT-3.5 and two well-trained natural language inference (NLI) models, Roberta-Large and XLNet-Large [31] from the huggingface hub [43], as the evaluators.
-With the same evaluators setup, the truthfulness evaluation follows the settings from [23], and uses the evaluation prompt provided in Appendix F.
+The fidelity follows the prior work Chen et al.
+(2025), which observes the discrepancy of the targeted LLMs by incorporating contrary hints to the input.
+The evaluation of truthfulness assesses the correlation between the derived NL explanations to the ground-truth explanations Specifically, we leverage GPT-4o and two well-trained natural language inference (NLI) models, Roberta-Large and XLNet-Large Nie et al.
+(2020) from the huggingface hub Wolf et al.
+(2019), as the evaluators.
+With the same evaluators setup, the truthfulness evaluation follows the settings from Liu et al.
+(2023), and uses the evaluation prompt provided in Appendix H.
 Specifically, the evaluators assess the derived explanations and ground-truth explanations, determining whether the two sentences belong to “similar content", “dissimilar content," or “non-relevant content".
-The higher proportion of “similar content", the more consistent results with the ground-truth NL explanations.
+Higher the proportion of “similar content", the more consistent results with ground-truth NL explanations.
 Implementation Details.
-In the experiments, we explore two variants of LLMs as the targeted LLMs $f(\cdot)$: Vicuna-7B [11] and Phi-2 [18], two types of LLMs as the explainers $\textsl{g}(\cdot)$ in FaithLM: GPT-3.5-Turbo and Claude-2 [3].
-The LLM agent for generating the contrary explanations takes the same LLMs as those used by the explainers.
-All reported results are calculated from the average scores of 3 times repetitions.
-The hyper-parameter settings and device configuration, including temperature and total optimization steps, of FaithLM are given in Appendix C and D, respectively.
-4.3 Explanation Efficacy of FaithLM (RQ1)
-In this section, we explore the efficacy of FaithLM from two perspectives.
-Specifically, we first evaluate the efficacy of NL explanations generated by Algorithm 1, and then assess the explanation trigger prompts generated by Algorithm 2.
-Efficacy of Derived Explanations.
-We assess the efficacy of derived explanations under the fidelity metric.
-FaithLM adopts the trajectory system prompts in Figure 13 of Appendix H.
-The generation of contrary explanations is guided by the prompt in Table 8 of Appendix G.
-Figure 3: The fidelity evaluation of explanations from FaithLM on ECQA dataset (left), TriviaQA-Long (middle), and COPA dataset (right).
-The reported scores are the average fidelity on testing instances in each step of fidelity-enhanced optimization.
-Figure 4: Trustfulness evaluation of the NL explanations.
-The higher the proportion of “similar to ground-truth explanation" obtained from classifiers, the more consistent the derived explanations are with the ground-truth NL explanations.
-  • Fidelity Evaluation.
-The results in Figure 3 demonstrate that FaithLM achieves significantly higher fidelity scores across all three datasets compared with two baselines after 20 steps of optimization.
-Moreover, the optimization curve of fidelity demonstrates that 20 rounds of optimization are sufficient to converge.
-Similar phenomenon occurs across different settings of explainers and targeted LLMs.
-Additional results are provided in Appendix E.
-  • Truthfulness Evaluation.
-To evaluate the truthfulness of the derived explanations, we show the proportions of “similar to ground-truth explanations" in the ECQA dataset, as depicted in Figure 4.
-We leverage GPT evaluator and well-trained NLI evaluators to determine whether the derived explanations from FaithLM and ground-truth explanations are within similar content or not.
-The results show that FaithLM’s derived explanations are more consistent with the ground-truth NL explanations, which is indicated by a larger proportion of “similar to ground-truth explanations" generated by FaithLM than baseline methods.
-Efficacy of Explanation Trigger Prompts.
-We first show prompt optimization curves of explanation trigger prompts on three different datasets, and then leverage the optimal explanation trigger prompts to generate explanations via FaithLM.
-In the experiments, we randomly select 15 instances from the training dataset in each optimization round, and compute the average fidelity scores of the newly derived trigger prompts.
-After the progress is terminated, we evaluate the optimized trigger prompts on the testing set.
-The optimization step is uniformly established at 50 rounds across different explainer and targeted LLM settings.
-We evaluate the derived explanation trigger prompts under two distinct experimental settings, as shown below.
-Figure 5: The fidelity in different optimization steps of the trigger prompts (Algorithm 2) on the ECQA, TrivaQA, and COPA datasets. The fidelity grows higher as the number of steps increases.
-  • Trigger Prompt Optimization Curve.
-Figure 5 demonstrates the optimization curves of three datasets.
-We display the explainer as GPT-3.5-Turbo and Claude-2 and the explainer as Vicuna-7B.
-We observe that the optimization curve exhibits a generally ascending trend as the step progresses, interspersed with multiple waves throughout the optimization procedure.
-This indicates that FaithLM generates better explanation trigger prompts after the optimization.
-More results of optimization curves on remaining datasets are provided in Appendix E.
-  • Explanation Generation by Optimized Trigger Prompts.
-We utilize the optimized explanation trigger prompts to generate explanations following Algorithm 1.
-The results are displayed in Figure 6(a), including the experiments conducted using all three datasets with Claude-2 as the explainer and Vicuna-7B as the targeted LLM.
-We observe that optimized explanation trigger prompts obtain higher fidelity scores than the initial human-crafted trigger prompt in generating explanations.
-This trend is consistent across all datasets, regardless of whether the explanations are refined by Algorithm 1.
-Additionally, the optimized trigger prompts are effective in improving the fidelity of the derived explanations with or without the optimization of FaithLM.
-A Case Study of  FaithLM.
-The case studies illustrate the evolving trend via FaithLM, including derived NL explanations, explanation trigger prompts, and contrary explanations in Appendix I.
-These studies demonstrate that the explanations generated by FaithLM are informative and readable, which enables humans to understand the reasons behind the decision-making process of LLMs.
-4.4 Transferability of Optimized Trigger Prompt (RQ2)
-We assess the transferability of ultimately optimized trigger prompts across different unseen datasets within the same domain, as depicted in Figure 6(b).
-Specifically, we transfer the optimized trigger prompts from the ECQA to the Social-IQA dataset, and from the COPA to the XCOPA datasets, without any additional optimization.
-Specifically, the Social-IQA dataset is dedicated to commonsense question-answering (similar to the ECQA dataset), while the XCOPA dataset specializes in causal reasoning (similar to the COPA dataset).
-We adopt the Vicuna-7B as the targeted LLM, and Claude-2 as the explainer on these transfer tasks.
-The fidelity of the derived NL explanation on the target dataset is shown in Figure 6(b).
-The optimized trigger prompts show better explanation efficacy than human-crafted prompts when it is transferred from a dataset from a similar domain.
-This shows that the optimized trigger prompts generated by FaithLM possess a great property of data transferability.
-(a) Robustness analytics of trigger prompts
-(b) Transferability results of trigger prompts
-Figure 6: Assessment on the adaptation of the optimized explanation trigger prompts. Figure (a) reveals the robustness evaluation, and figure (b) illustrates the results on transferability.
-(a) Evaluated by pre-trained NLI classifiers
-(b) Evaluated by GPT-3.5
-Figure 7: Ablation studies on evaluating contrary explanation. The results show that contrary explanations obtain opposite meanings to the derived explanations.
-4.5 Ablation Studies on Contrary Explanation (RQ3)
-The quality of contrary explanations $\neg\mathcal{E}_{\text{NL}}$ determines the efficacy of FaithLM.
-We leverage the powerful LLMs as the LLM agent to generate contrary explanations, requesting the delivery of high-quality opposite-meaning outputs from their original NL explanations.
-In this section, we evaluate the quality of contrary explanations, aiming to observe the semantic differences between the original NL explanations and their contrary explanations.
-To examine the quality, we employ the GPT classifier and two well-trained NLI classifiers, Roberta-Large and XLNet-Large [31].
-We leverage each classifier to distinguish whether the relationship between the “original NL explanations" and “contrary explanations" belong to the category of “similar meaning (entailment)," “dissimilar meaning (contradiction)," or “non-relevant (neutral)." We follow the evaluation settings from [23] on GPT-classifier with evaluation prompt provided in Table 7 of Appendix F.
-The results are shown in Figure 6 under the randomly sampled 100 instances from the ECQA and COPA datasets.
-We observe that the two NLI classifiers achieve up to 86% and 82% in the “dissimilar meanings"" category on the ECQA and COPA datasets, respectively.
-The empirical results of the GPT-classifier demonstrate that the derived explanations from SelfExp obtain more non-faithful information than FaithLM, which may increase the risk for the LLM agent of Fidelity Evaluator in generating non-relevant information as the contrary explanations.
-In addition, we provide case studies in Appendix I to demonstrate the informativeness and readability of contrary explanations.
-5 Conclusion
-In this paper, we introduce FaithLM to explain the decision-making process of LLMs.
-Specifically, FaithLM employs a fidelity enhancement strategy to progressively refine the fidelity of derived explanations and explanation trigger prompts.
-Experimental results on three datasets demonstrate the effectiveness of FaithLM in terms of the high fidelity of derived explanations, and better alignment with the ground-truth explanations of the datasets.
-As for future work, we plan to expand FaithLM to investigate high-stack applications, such as the healthcare domain, where transparency concerns are increasingly prominent due to the widespread adoption of black-box LLMs.
-Appendix
-Appendix A Details about Datasets
-The experiments are conducted on the three NLU datasets. The details of the datasets are provided as follows:
-  • ECQA [2].
-ECQA is an extension of the CQA dataset [37].
-Specifically, based on the CQA dataset, it annotates the positive or negative properties and golden explanations for the QA pairs.
-Due to API cost budgets, we evaluate our framework on the first 500 instances in the ECQA dataset.
-  • TriviaQA LongBench [19].
-TriviaQA LongBench (TriviaQA-Long) is a reading comprehension dataset.
-It includes 300 question-answer-evidence triples sourced from the Longbench [5] dataset [^1] https://huggingface.co/datasets/THUDM/LongBench/.
-This dataset features question-answer pairs crafted by trivia enthusiasts, accompanied by independently sourced evidence documents, providing supervision for answering these questions.
-  • Balanced COPA [35, 20].
-The Balanced COPA (COPA) dataset is a collection of 500 questions for commonsense causal reasoning.
-Each question consists of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
-Appendix B Related Work
-B.1 Post-hoc Explanation
+In the experiments, we explore two variants of LLMs as the targeted LLMs $f(\cdot)$: Vicuna-7B Chiang et al.
+(2023) and Phi-2 Javaheripi and Bubeck (2023), two types of LLMs as the explainers $\textsl{g}(\cdot)$ in FaithLM: GPT-3.5-Turbo and Claude-2 Anthropic (2023).
+The LLM agent for generating the contrary hints takes the same LLMs as those used by the explainers.
+All reported results are calculated from the average scores of 3 times repetitions with the grid search on the performance.
+The settings for predictors are uniform, with Phi-2 (2.7B) and Viucua-7B receiving identical hyperparameter configurations during the experiments conducted in this study.
+Appendix D Related Work
+D.1 Post-hoc Explanation
 Post-hoc explanation techniques have undergone significant development and discussion, driven by the widespread adoption of black-box ML models across various data modalities.
-A multitude of post-hoc algorithms has been introduced from two aspects: local and global explanations [30, 13].
-Explanations aim to explain the reasoning behind an individual model for each input instance, while global explanations aim to uncover the overall functioning of a complex model [12].
+A multitude of post-hoc algorithms has been introduced from two aspects: local and global explanations Molnar (2022); Du et al.
+(2019).
+Explanations aim to explain the reasoning behind an individual model for each input instance, while global explanations aim to uncover the overall functioning of a complex model Chuang et al.
+(2023).
 Considering various purposes of explanation, the explanation techniques mainly showcase the explanation from two perspectives, including feature attributions and counterfactual examples.
 Feature attribution aims to provide users with important scores for each feature’s impact on model predictions, while counterfactual examples aim to offer alternative instances that explicitly assist users in grasping the model’s decision-making process.
-In recent years, with the growing proficiency and wide usage of black-box LLMs, especially closed-source LLMs service, post-hoc explanations have become increasingly prominent and have garnered significant attention in NLP research due to the inaccessibility of LLMs’ model weights and structure [46].
-B.2 Explainability of LLMs
+In recent years, with the growing proficiency and wide usage of black-box LLMs, especially closed-source LLMs service, post-hoc explanations have become increasingly prominent and have garnered significant attention in NLP research due to the inaccessibility of LLMs’ model weights and structure Zhao et al.
+(2023).
+D.2 Explainability of LLMs
 The majority of explanation efforts in LLM research have centered on delivering explanations.
-One group of studies calculates importance scores for specific tokens [24, 17], another line of progress generates NL explanations by leveraging the pre-trained LLMs with internal model knowledge sources [21, 9, 28], the other group of work leverages LLMs themselves to generate chain-of-thought (CoT) reasoning [22, 34, 8, 8] as the self-explanations through the one feed-forward inference process.
-Furthermore, some studies aim to yield counterfactual explanations by pre-trained LLMs to assist users in better understanding the decision-making process from LLMs [7, 8].
+One group of studies calculates importance scores for specific tokens Lopardo et al.
+(2023); Huang et al.
+(2023), another line of progress generates NL explanations by leveraging the pre-trained LLMs with internal model knowledge sources Kumar and Talukdar (2020); Chen et al.
+(2023c); Menon et al.
+(2023), the other group of work leverages LLMs themselves to generate chain-of-thought (CoT) reasoning Lanham et al.
+(2023); Radhakrishnan et al.
+(2023); Chen et al.
+(2023b, b) as the self-explanations through the one feed-forward inference process.
+Furthermore, some studies aim to yield counterfactual explanations by pre-trained LLMs to assist users in better understanding the decision-making process from LLMs Chen et al.
+(2021, 2023b).
 Although NL explanations offer fantastic human-understandable insights than token-wise explanations, the explanations can lose their fidelity via one feed-forward inference process of pre-trained LLMs.
-Unreliability and non-fidelity of NL explanations are still a concern [45, 40].
-Given our primary aim of producing faithful explanations, our efforts are concentrated on generating NL explanations to improve the likelihood of accurately representing the decision-making process of LLMs.
-B.3 LLMs as Optimizers
-LLMs as optimizers is a novel paradigm, describing optimization problems in natural language and utilizing the reasoning capabilities of LLMs for optimizing [44].
-Depicting optimization problems in natural language enables the optimization of diverse tasks without defining formal specifications, such as prompt optimization [44, 10, 16], agent learning [36], and model labeling [38].
+Unreliability and non-fidelity of NL explanations are still a concern Ye and Durrett (2022); Turpin et al.
+(2023).
+Given our primary aim of producing faithful explanations, our efforts are to generate NL explanations to improve the likelihood of accurately representing the decision-making process of LLMs.
+D.3 LLMs as Optimizers
+LLMs as optimizers is a novel paradigm, describing optimization problems in natural language and utilizing the reasoning capabilities of LLMs for optimizing Yang et al.
+(2023).
+Depicting optimization problems in natural language enables the optimization of diverse tasks without defining formal specifications, such as prompt optimization Yang et al.
+(2023); Cheng et al.
+(2023); Guo et al.
+(2023), agent learning Shinn et al.
+(2023), and model labeling Thomas et al.
+(2023).
 Based on this optimization paradigm, our work introduces a generative explanation framework with a novel estimation method of sentence-level fidelity.
-Appendix C Hyper-parameter Settings of FaithLM
+Appendix E Hyper-parameter Settings of FaithLM
 The hyper-parameters of FaithLM are given in Table 2.
 The configuration for explainers is consistent across Claude-2 and GPT-3.5-Turbo, provided that the parameters are adjustable.
-Likewise, the settings for predictors are uniform, with Phi-2 and Viucua-13B receiving identical hyperparameter configurations during the experiments conducted in this study.
+Likewise, the settings for predictors are uniform, with Phi-2 and Viucua-7B receiving identical hyperparameter configurations during the experiments conducted in this study.
 Dataset ECQA TriviaQA-Long COPA
-Fidelity-enhanced Optimization Optimization Steps 20 20 20
+Fidelity-enhanced
+Optimization
+Optimization Steps 20 20 20
 Temperature of Predictor LLMs 0.7 0.5 0.7
@@
 Top-P of Explainer LLMs 0.9 0.9 0.9
-Trigger-oriented Optimization Optimization Steps 50 100 100
+Trigger-oriented
+Optimization
+Optimization Steps 50 100 100
 Sampled Instances 30 30 30
@@
 Table 2: Hyper-parameters and optimization settings in FaithLM.
-Appendix D Computation Infrastructure and Costs
-D.1 Computation Infrastructure
+Appendix F Computation Infrastructure and Costs
+F.1 Computation Infrastructure
 For a fair comparison of testing algorithmic throughput, the experiments are conducted based on the following physical computing infrastructure in Table 3.
@@
 Table 3: Computing infrastructure for the experiments.
-D.2 Computation Costs
-The computational costs associated with xLLM primarily differ from the inference costs of local LLMs and the expenses related to API-accessed LLMs.
-The computational costs depend on the parameter scale and variants of LLMs used in the xLLM framework, shown in Table 4 and 5.
+F.2 Computation Costs
+The computational costs associated with FaithLM primarily differ from the inference costs of local LLMs and the expenses related to API-accessed LLMs.
+The computational costs depend on the parameter scale and variants of LLMs used in the FaithLM framework, shown in Table 4 and 5.
 ECQA TrivaQA COPA
@@
 Table 5: Computing costs of FaithLM with Vicuna-7B under different batch size (bs).
-Appendix E Additional Experimental results of FaithLM
-E.1 Optimization Procedure of derived explanations
+Appendix G Additional Experimental results of FaithLM
+G.1 Optimization Procedure of derived explanations
 We demonstrate more evaluation results on derived explanations from FaithLM.
@@
 Figure 8: The fidelity evaluation of derived explanations from FaithLM under different settings of predictors and explainers.
-E.2 Robustness Analytics of Configurations in FaithLM
+G.2 Additional Optimization Curve of Explanation Trigger Prompt
+We demonstrate more evaluation results on the optimization curve of explanation trigger prompts of FaithLM.
+The optimization curve shown in Figure 9 generally displays an upward trend with the progression of steps, interspersed with several fluctuations throughout the optimization process.
+This suggests that FaithLM can successfully generate improved explanation trigger prompts after optimization.
+Figure 9: The optimization curve of explanation trigger prompts on ECQA (left), TriviaQA-Long (middle), and COPA dataset (right).
+G.3 Additional Experiments on Diverse Domains of Dataset
+We further have conducted additional experiments on one new MedMCQA dataset Pal et al.
+(2022) in the healthcare domain.
+We evaluate the FaithLM using a fidelity assessment under Natural Language Explanation Generation settings.
+All experimental configurations follow the settings in Section 4.2.
+The experimental results are shown in the table below.
+We observe that FaithLM outperforms the baseline method, which is consistent with the experimental results across other domain datasets that were evaluated in our work.
+SelfExp Self-consistency FaithLM
+Fidelity 0.6956 0.4715 0.9565
+Table 6: Additional experimental results on MedMCQA dataset.
+G.4 Additional Experiments on Truthfulness Evaluation
+We evaluate all baseline methods and FaithLM under multiple settings to assess how closely the derived explanations match the ground-truth rationales.
+A GPT-based evaluator assigns a GPT-Score from 1 to 5, with higher values indicating greater semantic similarity.
+Explanations tagged as “similar content” or scoring near 5 are treated as matches.
+We report the truthfulness of generated explanations in Figure 10.
+A GPT-based evaluator assigns a truthfulness score from 1 to 5 by checking factual consistency with the input, the task label, and commonsense knowledge.
+Explanations tagged as factually consistent or scoring near 5 are counted as truthful.
+Across all settings, FaithLM attains the highest mean truthfulness score and the largest fraction of truthful explanations, indicating that our optimization improves not only fidelity but also factual quality of the produced rationales.
+Figure 10: Truthfulness evaluation with ground-truth explanation under GPT-score settings.
+G.5 Additional Experiments of Contrary Hints.
+We here showcase the results of using NLI classifiers to evaluate the quality of contrary hints.
+Figure 11: Quality Evaluation results of Contrary Hints via NLI classifiers
+Appendix H Details of Evaluation Prompt Usage
+We provide a listing of the evaluation prompts in Table 7 utilized in assessing the performance of FaithLM.
+The first row reveals the evaluation prompt on comparing the derived explanation with the ground-truth (GT) explanation in the ECQA dataset in Section 4.3; and the second row demonstrates the evaluation prompt on activating the GPT classifier and the GPT scorer for assessing contrary hints in Section 4.5.
+Evaluation Task Evaluation Prompts
+Ground-truth Explanation Given a user instruction and two AI assistant responses, your job is to classify whether the relation of two responses in S1 and S2 belongs to G-1, G-2, or G-3.
+The meaning of class is as follows: (G-1) relevant contents, (G-2) irrelevant contents, or (G-3) irrelevant contents.
+Judge responses holistically, paying special attention to whether two responses have similar contents.
+Judge responses with only ONE class label as your final answer.
+S1:{derived explanation}.
+S2:{GT-Explanation}.
+Please ONLY response your in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
+GPT classifier for contrary hints Given a user instruction and two AI assistant responses, your job is to classify whether two responses in S1 and S2 belong to G-1, G-2, or G-3.
+The meaning of class is as follows: (G-1) same semantic meaning, (G-2) opposite semantic meaning, and (G-3) no relation.
+Judge responses holistically, paying special attention to whether two responses have the same semantic meaning.
+Judge responses with only ONE class label as your final answer.
+S1:{derived explanation}.
+S2:{contrary hints}.
+Please ONLY respond in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
+GPT scorer of contrary hints Given a user instruction and two AI assistant responses, your job is to rate from ONE to FIVE to judge whether two responses in S1 and S2 have the same semantic meaning or not.
+A FIVE score refers to being totally the same, and ONE score refers to being totally the opposite.
+Judge responses holistically, paying special attention to whether two responses have the same semantic meaning.
+The judge responds with the rates between ONE and FIVE.
+S1:{derived explanation}.
+S2:{contrary hints}.
+Please ONLY respond to the rate value; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
+Table 7: Evaluation Prompts given to GPT-3.5-Turbo used in assessing the efficacy of FaithLM.
+H.1 Robustness Analytics of Configuration (RQ3)
 In this section, the robustness test of the explainer LLMs is conducted under the analytics of hyper-parameters that are highly dependent on the outputs of LLMs.
@@
 Thus, in the main experiments, all reported performances are under the settings of temperature 0.9 and top-p 0.9, achieving the best performance for generating explanations.
-Figure 9: Robustness Analytics of FaithLM on different LLM settings, which are Temperature (left) and Top-p strategies (right).
-E.3 Additional Optimization Curve of Explanation Trigger Prompt
-We demonstrate more evaluation results on the optimization curve of explanation trigger prompts of FaithLM.
-The optimization curve shown in Figure 10 generally displays an upward trend with the progression of steps, interspersed with several fluctuations throughout the optimization process.
-This suggests that FaithLM can successfully generate improved explanation trigger prompts after optimization.
-Figure 10: The optimization curve of explanation trigger prompts on ECQA (left), TriviaQA-Long (middle), and COPA dataset (right).
-E.4 Additional Experiments on Diverse Domains of Dataset
-We further have conducted additional experiments on one new MedMCQA dataset [32] in the healthcare domain.
-We evaluate the FaithLM using a fidelity assessment under Natural Language Explanation Generation settings.
-All experimental configurations follow the settings in Section 4.2.
-The experimental results are shown in the table below.
-We observe that FaithLM outperforms the baseline method, which is consistent with the experimental results across other domain datasets that were evaluated in our work.
-SelfExp Self-consistency FaithLM
-Fidelity 0.6956 0.4715 0.9565
-Table 6: Additional experimental results on MedMCQA dataset.
-E.5 Additional Experiments on Truthfulness Evaluation
-We further lunch additional experiments on all evaluators with different evaluation settings to measure the relevance between derived explanations and ground-truth explanations.
-The evaluators assesses their relevance using a GPT-Score in GPT evaluator on a scale of one to five.
-If the two provided explanations are classified under the “similar content" category or receive a GPT-Score close to five, this indicates that the derived explanations are highly similar to ground-truth explanations.
-Figure 11: Truthfulness evaluation with ground-truth explanation under GPT-score settings.
-Appendix F Details of Evaluation Prompt Usage
-We provide a listing of the evaluation prompts in Table 7 utilized in assessing the performance of FaithLM.
-The first row reveals the evaluation prompt on comparing the derived explanation with the ground-truth (GT) explanation in the ECQA dataset in Section 4.3; and the second row demonstrates the evaluation prompt on activating the GPT classifier and the GPT scorer for assessing contrary explanations in Section 4.5.
-Evaluation Task Evaluation Prompts
-Ground-truth Explanation Given a user instruction and two AI assistant responses, your job is to classify whether the relation of two responses in S1 and S2 belongs to G-1, G-2, or G-3.
-The meaning of class is as follows: (G-1) relevant contents, (G-2) irrelevant contents, or (G-3) irrelevant contents.
-Judge responses holistically, paying special attention to whether two responses have similar contents.
-Judge responses with only ONE class label as your final answer.
-S1:{derived explanation}.
-S2:{GT-Explanation}.
-Please ONLY response your in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
-GPT classifier for Contrary Explanations Given a user instruction and two AI assistant responses, your job is to classify whether two responses in S1 and S2 belong to G-1, G-2, or G-3.
-The meaning of class is as follows: (G-1) same semantic meaning, (G-2) opposite semantic meaning, and (G-3) no relation.
-Judge responses holistically, paying special attention to whether two responses have the same semantic meaning.
-Judge responses with only ONE class label as your final answer.
-S1:{derived explanation}.
-S2:{Contrary Explanations}.
-Please ONLY respond in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
-GPT scorer of Contrary Explanations Given a user instruction and two AI assistant responses, your job is to rate from ONE to FIVE to judge whether two responses in S1 and S2 have the same semantic meaning or not.
-A FIVE score refers to being totally the same, and ONE score refers to being totally the opposite.
-Judge responses holistically, paying special attention to whether two responses have the same semantic meaning.
-The judge responds with the rates between ONE and FIVE.
-S1:{derived explanation}.
-S2:{Contrary Explanations}.
-Please ONLY respond to the rate value; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
-Table 7: Evaluation Prompts given to GPT-3.5-Turbo used in assessing the efficacy of FaithLM.
-Appendix G Details of Prompts Usage in FaithLM
+Figure 12: Robustness Analytics of FaithLM: Temperature (left) and Top-p strategies (right).
+Appendix I Details of Prompts Usage in FaithLM
 We provide a listing of the prompts in Table 7 utilized in FaithLM in different tasks.
 The first row demonstrates the initial explanation trigger prompt leveraging in both fidelity-enhanced optimization and trigger-oriented optimization.
-The second row shows the prompt for the LLM agent to generate the contrary explanations.
+The second row shows the prompt for the LLM agent to generate the contrary hints.
 Conducted Task Evaluation Prompts
@@
 A:{Targeted LLM-generated Answer}.
-Contrary Explanations Generation Please generate one example of obtaining the opposite meaning from a given sentence. Make sure you output sentences only. Sentences:{derived explanation}.
-Table 8: The example of prompts that are given to two explainer LLMs and LLM agent for contrary explanation.
-Appendix H Trajectory System Prompts Usage in FaithLM
+contrary hints Generation Please generate one example of obtaining the opposite meaning from a given sentence. Make sure you output sentences only. Sentences:{derived explanation}.
+Table 8: The example of prompts that are given to two explainer LLMs and LLM agent for contrary hint.
+Appendix J Trajectory System Prompts Usage in FaithLM
 We present a detailed listing of the trigger-oriented trajectory prompt in Figure 13 and the explanation-oriented prompt in Figure 14, as utilized within the FaithLM framework.
-H.1 Trigger-oriented Trajectory Prompt
+J.1 Trigger-oriented Trajectory Prompt
 […]
@@
 The output of FaithLM optimized under trigger-oriented trajectory prompt is append after the Response label.
-H.2 NL Explanation-oriented Trajectory Prompt
+J.2 NL Explanation-oriented Trajectory Prompt
 […]
@@
 The output of FaithLM optimized under trigger-oriented trajectory prompt is append after the Response label.
-Appendix I Case Studies of FaithLM
-I.1 Explanations Generation
+Appendix K Case Studies of FaithLM
+K.1 Explanations Generation
 We demonstrate the case studies of explanation results generated by FaithLM after fidelity-enhanced optimization in the following two Figure 15 and Figure 16.
@@
 Figure 16: Case Study: Derived explanations (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Phi)
-I.2 Explanation Trigger Prompts
+K.2 Explanation Trigger Prompts
 The demonstrations in the explanation trigger prompts generated by FaithLM in Figure 17.
@@
 Figure 17: Case Study: Generation of Explanation Trigger Prompts
-I.3 Contrary Explanation
-We demonstrate results on contrary explanations leveraged in Fidelity Evaluator. The results in Figure 17 show that the contrary explanation obtains the opposite meaning of the given explanation.
+K.3 contrary hint
+We demonstrate results on contrary hints leveraged in Fidelity Evaluator. The results in Figure 18 show that the contrary hint obtains the opposite meaning of the given explanation.
 Case Study 1 Question: The child wanted to test electricity, so what did he feel as a result?
@@
 It failed to comprehend the potential outcomes of unsafe electrical contact.
-Contrary Explanation: The model correctly predicted that testing electricity would not cause a fire.
+contrary hint: The model correctly predicted that testing electricity would not cause a fire.
 It likely fully understood that direct contact with electricity can cause a shock rather than ignite a fire.
@@
 It has compartments to store small items like coins, so putting coins in your purse allows you to easily bring them along wherever you go.
-Contrary Explanation: The purse is not a good place to put coins if you don’t want to bring them with you, because purses are designed for other items, not coins.
-Figure 17: Case Study: Contrary Explanation (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)
-I.4 Limitations of Learning FaithLM
-One significant limitation of FaithLM associated with the carbon emissions during the experiments.
-The extensive computational power required for optimizing FaithLM leads to considerable energy consumption, which, in turn, results in a significant carbon footprint.
-As the demand for more sophisticated LLMs continues to grow, so does their environmental impact.
-This limitation underscores the urgent need to explore and adopt more sustainable practices and technologies in the development and learning FaithLM with fewer optimization steps to mitigate their ecological footprint.
+contrary hint: The purse is not a good place to put coins if you don’t want to bring them with you, because purses are designed for other items, not coins.
+Figure 18: Case Study: contrary hint (Dataset: ECQA; Targeted Model $f(\cdot)$: Claude-2; Explainer $\textsl{g}(\cdot)$: Vicuna)